### v1

# Modified Langevin noise formalism for multiple quantum emitters in dispersive electromagnetic environments

Abstract
The control of interactions among quantum emitters through nanophotonic structures offers significant potential for quantum technologies. However, a theoretical description of the interaction of multiple quantum emitters with complex dispersive dielectric objects remains highly challenging. Here we introduce an approach based on the modified Langevin noise formalism that unveils the roles of both the noise polarization currents of the dielectrics and the vacuum fluctuations of the electromagnetic field scattered by the dielectrics. This extends Refs. [1], [2] to the general case of an arbitrary number of emitters. The proposed approach allows us to describe the dynamics of the quantum emitters for arbitrary initial quantum states of the electromagnetic environment consisting of two independent bosonic reservoirs, a medium-assisted reservoir and a scattering-assisted reservoir, each characterized by its own spectral density matrix. Understanding how these reservoirs shape emitter dynamics is crucial to understanding light–matter interactions in complex electromagnetic environments and to enhancing intrinsic emitter properties in structured environments.
I Introduction
The design of qubit-qubit interactions is fundamental to a wide range of quantum technologies, including quantum networking, quantum information processing, and quantum computation (e.g., [3], [4], [5], [6]). Nanophotonics enables the engineering and control of quantum properties of light by embedding quantum emitters within nanostructured environments. In particular, systems that utilize quantum emitters as qubits have attracted growing interest [7, 8], with significant efforts devoted to tailoring their mutual interactions through metallic or dielectric nano-structures [9]. These systems are physically rich: metal and dielectric nanoparticles, typically dispersive and lossy, can support plasmonic and dielectric resonances of various multipolar orders, which can be engineered by adjusting the shape and spatial arrangement of the nanoparticles, or by exploiting collective resonances. Moreover, nanophotonic devices enable subwavelength confinement of light, allowing device miniaturization, enhanced light–matter interactions, and consequently faster dynamics and higher operating speeds. Such engineered couplings play a central role in the generation and manipulation of non-classical states of light. Nevertheless, providing a theoretical description of multiple quantum emitters that interact with realistic nanostructures remains highly nontrivial. The electromagnetic environment is inherently dispersive, and its modes span a high-dimensional continuum. The account of the finite extent, material losses, and dispersion of the dielectric components poses serious challenges to the canonical quantization of the electromagnetic field. In this context, macroscopic quantum electrodynamics (e.g. [10], [11]), offers a widely adopted phenomenological framework, enabling a consistent and flexible quantization scheme in the presence of complex dielectric media.
Macroscopic quantum electrodynamics relies on the Langevin noise formalism, which is based on the fluctuation dissipation theorem. The electromagnetic field, commonly referred to as the medium-assisted field, emerges from the dielectric noise polarization current, mediated by the dyadic Green function of the dielectric objects. As argued in refs. [12] and [13] the original macroscopic quantum electrodynamic model disregards the influence of vacuum fluctuations of the electromagnetic field scattered by the dielectric objects, called the scattering-assisted field in [14] and [15]. The modified Langevin noise formalism adds the scattering-assisted field to the medium-assisted field: polarization current fluctuations and vacuum fluctuations of the electromagnetic field scattered by the dielectric objects are on the same footing. Recently, it has been justified for finite-size dispersive dielectric objects with arbitrary shapes applying the Heisenberg picture to a phenomenological model of dielectric media based on a continuum set of harmonic oscillators [15]. Each elementary region of the dielectric medium is described as a bath of harmonic oscillators that couples to the electromagnetic field in such a way that the quantized electromagnetic field experiences the dielectric medium through its macroscopic dielectric permittivity (e.g., [16], [17]). As the medium-assisted field, the scattering-assisted field can also be expressed in terms of vacuum fluctuations through the dyadic Green function of the dielectric objects [15].
The modified Langevin noise formalism has recently been applied for the first time to a single quantum emitter that interacts with a linear dispersive dielectric slab in [14], [1], and with a linear dispersive dielectric sphere in [2]. The quantum emitter interacts with two distinct and independent bosonic reservoirs, the medium-assisted reservoir and the scattering-assisted reservoir, each characterized by its own spectral density. The medium-assisted and the scattering-assisted spectral densities are quadratic functionals of the dyadic Green function of dielectric objects $\mathcal{G}_{\omega}(\mathbf{r},\mathbf{r}^{\prime})$. The time evolution of the reduced density operator of the quantum emitter depends on both the initial quantum state of the entire system and the spectral densities of the two reservoirs. Only under specific conditions can the actions of the two reservoirs combine in such a way that they can be represented by a single equivalent bosonic reservoir characterized by an equivalent spectral density depending only on the value of the dyadic Green function of the dielectric object at the quantum emitter position. In fact, when the initial quantum state of the entire system is a product state and both reservoirs are initially in the vacuum state, the reduced-density operator of the quantum emitter evolves as the emitter interacts with an equivalent single bosonic reservoir, with only positive frequencies and equivalent spectral density $\mathcal{J}(\omega)$ equal to the sum of the spectral density of the individual reservoirs. In this case, it results that [14], [1], [2], $\mathcal{J}(\omega)=\frac{\omega^{2}}{\pi\hbar\varepsilon_{0}c^{2}}\mathbb{p}\cdot\operatorname{Im}\left[\mathcal{G}_{\omega}\left(\mathbf{r}_{a},\mathbf{r}_{a}\right)\right]\cdot\mathbb{p}$ as a consequence of the fundamental integral relation 31 of ref. [15]; here $\mathbf{r}_{a}$ is the position vector of the emitter and $\mathbb{p}$ is its transition dipole moment. When the two reservoirs are initially in thermal quantum states with equal temperature $T_{0}$ the reduced dynamics of the quantum emitter can yet be described by a single zero-temperature bosonic reservoir with positive and negative frequencies and an equivalent temperature-dependent spectral density equal to $\theta(\omega,T_{0})\mathcal{J}(|\omega|)$ where $\theta(\omega,T_{0})=\frac{1}{2}sign(\omega)\left[1+\coth\left(\frac{\beta_{0}\hbar\omega}{2}\right)\right]$ and $\beta_{0}=1/(k_{B}T_{0})$. In the literature based on the Langevin noise formalism, similar relations are widely used; however, the conditions under which they remain valid are not always clearly stated. Indeed, if the two reservoirs have different temperatures, simply knowing the values of the dyadic Green function of the dielectric objects at the position of the quantum emitter is not sufficient to evaluate the dynamics of the emitter. In this case, the reduced dynamics of the emitter can yet be described by a zero temperature single bosonic reservoir with an equivalent spectral density equal to the sum of the temperature-dependent medium-assisted spectral density and the temperature-dependent scattering-assisted spectral density [2]. In general, if the initial states of the two reservoirs are not thermal states, the situation becomes considerably more complicated, and it is no longer possible to describe their action through an equivalent single reservoir.
Although the model introduced in refs. [1], [2] is capable of handling an electromagnetic environment with arbitrary initial quantum state, the considered framework is limited to a single quantum emitter. In the present article, we lift this restriction and generalize the approach based on the modified Langevin noise formalism to a collection of multiple quantum emitters with arbitrary orientations of the transition dipole moments. As we did for the single quantum emitter [1], we applied the emitter-centered mode technique to reduce the number of degrees of freedom of the electromagnetic environment (e.g. [18], [19]). In particular, we show also in this case that the reduced dynamic of the quantum emitters depends only on the values of the dyadic Green function at the quantum emitter positions if the initial state of the entire system is a product state and the initial states of the medium-assisted and the scattering-assisted reservoirs are thermal states with the same temperature. When these conditions are not satisfied, the reduced dynamics of the quantum emitter depends on the generalized spectral densities of the single reservoirs, which are quadratic functionals of the dyadic Green function of the dielectric objects. We first introduce the generalized spectral density matrices that characterize the medium-assisted and the scattering-assisted reservoirs and discuss their general properties. Then, we introduce an equivalent reduced electromagnetic environment characterized by a single spectral density matrix for scenarios in which the two bosonic reservoirs are initially in thermal quantum states with different temperatures. We eventually applied the approach to study the correlation between two quantum emitters interacting with a Drude sphere or with a Drude nanostructure composed of two rods and a disk. We first compute the medium-assisted and the scattering-assisted spectral density matrices for the two cases, then we evaluate the evolution of the reduced density operator of the two quantum emitters and find entanglement decay and revivals, as well as entanglement generation.
In Section II we apply the modified Langevin noise formalism to model multiple quantum emitters interacting with a dispersive electromagnetic environment. In Section III, we use emitter-centered modes to reduce the number of degrees of freedom of the electromagnetic environment. In Section IV we introduce the medium- and the scattering-assisted spectral density matrices. In Section V, we propose a surrogate bosonic environment to describe the reduced dynamics of quantum emitters when the initial quantum state of the entire system is a product state, and the medium-assisted and scattering-assisted bosonic reservoirs are initially in thermal quantum states with different temperatures. In Section VI we apply our model to a system of two quantum emitters. In Section VII, we give a summary and discuss our conclusions.
II Model
A collection of $N$ quantum emitters mutually interacts and couples to finite-size dielectric objects of arbitrary shape embedded in an unbounded space. We denote by $\mathbf{r}_{j}$ the position vector of the $j$-th quantum emitter with j=1,2,,Nj=1,2,...,N. We assume that the dielectrics are linear, isotropic, and dispersive in time. We denote by $V$ the region occupied by the dielectric objects and by $\varepsilon_{\omega}(\mathbf{r})$ their relative dielectric permittivity in the frequency domain. The dispersive dielectrics together with the electromagnetic field constitute the electromagnetic environment of the quantum emitters.
II.1 Hamiltonian
The Hamiltonian of the entire system, multiple quantum emitters ++ electromagnetic environment, is given by
$\hat{H}=\hat{H}_{A}+\hat{H}_{E}+\hat{H}_{I},$ (1)
where $\hat{H}_{A}$ is the bare Hamiltonian of the quantum emitters, assumed to be mutually isolated, $\hat{H}_{E}$ is the bare Hamiltonian of the electromagnetic environment, and $\hat{H}_{I}$ describes their interaction.
In the multipolar coupling scheme (Power–Zienau–Woolley picture) and within the dipole approximation the interaction Hamiltonian $\hat{H}_{I}$ is given by
$\hat{H}_{I}=-\sum_{j=1}^{N}\hat{\mathbb{d}}_{j}\cdot\hat{\mathbf{E}}(\mathbf{r}_{j})$ (2)
where $\hat{\mathbf{E}}(\mathbf{r})$ is the electric field operator and $\hat{\mathbb{d}}_{j}$ is the electric dipole moment operator of the $j$-th emitter located at $\mathbf{r}_{j}$. We assume that the dipole moment operator of the $j$-th emitter couples through a fixed polarization direction $\mathbb{u}_{j}$ as in ref. [19], so that $\hat{\mathbb{d}}_{j}=\hat{\mu}_{j}\mathbb{u}_{j}$ where $\hat{\mu}_{j}$ is the corresponding transition dipole moment operator. More generally, one may retain up to three orthogonal polarization components per emitter, $\hat{\mathbb{d}}_{j}=\sum_{a}\hat{\mu}_{ja}\mathbb{a}$ where $a=x,y,z$, and $\hat{\mu}_{jx},\hat{\mu}_{jy},\hat{\mu}_{jz}$ are the corresponding transition-dipole moment operators.
II.2 Diagonal Form of the Electromagnetic Environment Hamiltonian
The bare Hamiltonian of the electromagnetic environment $\hat{H}_{E}$ accounts for the electromagnetic field, the polarization currents of the dielectric objects, and their interaction. The modified Langevin noise formalism provides a straightforward way to diagonalize it [15].
We express the electric field operator in the Schrödinger picture as $\hat{\mathbf{E}}(\mathbf{r})=\int^{\infty}_{0}d\omega[\hat{\mathbf{E}}_{\omega}(\mathbf{r})+h.c.]$ where $\hat{\mathbf{E}}_{\omega}(\mathbf{r})$ is its monochromatic component in the Heisenberg picture. In the modified Langevin noise formalism, the electric field operator $\hat{\mathbf{E}}_{\omega}(\mathbf{r})$ has two contributions, the medium-assisted contribution $\hat{\mathbf{E}}^{(M)}_{\omega}(\mathbf{r})$ and the scattering-assisted contribution $\hat{\mathbf{E}}^{(S)}_{\omega}(\mathbf{r})$,
$\hat{\mathbf{E}}_{\omega}=\hat{\mathbf{E}}_{\omega}^{(M)}+\hat{\mathbf{E}}_{\omega}^{(S)}.$ (3)
The medium-assisted contribution is generated by the noise polarization currents of the dielectric object [10], while the scattering-assisted contribution is generated by the vacuum fluctuations of the electromagnetic field scattered by the dielectric object [12], [13]. These two contributions are expressed in terms of particular bosonic operators that diagonalize $\hat{H}_{E}$.
The medium-assisted field $\hat{\mathbf{E}}_{\omega}^{(M)}(\mathbf{r})$ is given by [15]
$\hat{\mathbf{E}}_{\omega}^{(M)}(\mathbf{r})=\int_{V}d^{3}\mathbf{r}^{\prime}\,\mathcal{G}_{e}\left(\mathbf{r},\mathbf{r}^{\prime};\omega\right)\hat{\mathbf{f}}_{\omega}\left(\mathbf{r}^{\prime}\right),$ (4)
where $\hat{\mathbf{f}}_{\omega}\left(\mathbf{r}\right)$ is the monochromatic bosonic field operator describing the fluctuations of the dielectric polarization currents, whose support is the region $V$,
$\mathcal{G}_{e}\left(\mathbf{r},\mathbf{r}^{\prime};\omega\right)=i\frac{\omega^{2}}{c^{2}}\sqrt{\frac{\hbar}{\pi\varepsilon_{0}}\operatorname{Im}\left[\varepsilon_{\omega}\left(\mathbf{r}^{\prime}\right)\right]}\mathcal{G}_{\omega}\left(\mathbf{r},\mathbf{r}^{\prime}\right),$ (5)
$\mathcal{G}_{\omega}\left(\mathbf{r},\mathbf{r}^{\prime}\right)$ is the dyadic Green’s function of the dielectric object satisfying the equation
$\left(\nabla_{\mathbf{r}}\times\nabla_{\mathbf{r}}\times-k_{\omega}^{2}\varepsilon_{\omega}\right)\mathcal{G}_{\omega}\left(\mathbf{r},\mathbf{r}^{\prime}\right)=\delta\left(\mathbf{r}-\mathbf{r}^{\prime}\right)\mathcal{I},$ (6)
and the boundary condition $\mathcal{G}_{\omega}\left(\mathbf{r},\mathbf{r}^{\prime}\right)\rightarrow 0$ for $r,r^{\prime}\rightarrow$ $\infty$; $\varepsilon_{0}$ is the vacuum permittivity, $k_{\omega}=\omega/c$, $c$ is the light velocity in vacuum, and $\mathcal{I}$ is the identity dyad.
Let $\mathbf{F}_{\omega\mathbf{n}\nu}(\mathbf{r})$ denote the solution of homogeneous equation
$\left(\nabla_{\mathbf{r}}\times\nabla_{\mathbf{r}}\times-k_{\omega}^{2}\varepsilon_{\omega}\right)\mathbf{F}_{\omega\mathbf{n}\nu}=0$ (7)
obeying the boundary condition
$\mathbf{F}_{\omega\mathbf{n}\nu}(\mathbf{r})\underset{r\rightarrow\infty}{\approx}e^{ik_{\omega}\mathbf{r}\cdot\mathbf{n}}\mathbb{e}_{\mathbf{n}\nu},$ (8)
where $\mathbf{n}$ is the unit vector along the wave vector $\mathbf{k}=k_{\omega}\mathbf{n}$ and $\mathbf{e}_{\mathbf{n}1}$, $\mathbf{e}_{\mathbf{n}2}$ are two mutually orthogonal polarization unit vectors orthogonal to $\mathbf{n}$. We introduce the scattering modes $\mathbf{E}_{\omega\mathbf{n}\nu}(\mathbf{r})$ defined as [15]
$\mathbf{E}_{\omega\mathbf{n}\nu}(\mathbf{r})=\sqrt{\frac{\hbar\omega^{3}}{16\pi^{3}\varepsilon_{0}c^{3}}}\mathbf{F}_{\omega\mathbf{n}\nu}(\mathbf{r}).$ (9)
The scattering-assisted field $\hat{\mathbf{E}}_{\omega}^{(S)}$ is given by
$\hat{\mathbf{E}}^{(S)}_{\omega}(\mathbf{r})=\oint do_{\mathbf{n}}\sum_{\nu}\mathbf{E}_{\omega\mathbf{n}\nu}(\mathbf{r})\hat{g}_{\omega\mathbf{n}\nu},$ (10)
where $\hat{g}_{\omega\mathbf{n}\nu}$ is the monochromatic bosonic operator that describes the fluctuation of the radiation incoming from infinity and scattered by the dielectric object. Here $o_{\mathbf{n}}=(\theta_{\mathbf{n}},\phi_{\mathbf{n}})$ are the polar angles of the unit vector $\mathbf{n}$}, $do_{\mathbf{n}}=\sin\theta_{\mathbf{n}}d\theta_{\mathbf{n}}d\phi_{\mathbf{n}}$ is the differential of the solid angle and the integration is performed over the whole solid angle with $\theta\in[0,\pi]$ and $\phi\in[0,2\pi]$.
The bosonic field operators $\hat{\mathbf{f}}_{\omega}(\mathbf{r})$ and $\hat{g}_{\omega\mathbf{n}\nu}$ are independent. Any possible commutation relations between them vanishes except the fundamental ones,
${\left[\hat{\mathbf{f}}_{\omega}(\mathbf{r}),\hat{\mathbf{f}}_{\omega^{\prime}}^{\dagger}\left(\mathbf{r}^{\prime}\right)\right]}$ $=\delta\left(\omega-\omega^{\prime}\right)\delta\left(\mathbf{r}-\mathbf{r}^{\prime}\right)I,$ (11a)
${\left[\hat{g}_{\omega\mathbf{n}\nu},\hat{g}_{\omega^{\prime}\mathbf{n}^{\prime}\nu^{\prime}}^{\dagger}\right]}$ $=\delta\left(\omega-\omega^{\prime}\right)\delta\left(o_{\mathbf{n}}-o_{\mathbf{n}^{\prime}}\right)\delta_{\nu\nu^{\prime}},$ (11b)
where $\delta\left(o_{\mathbf{n}}-o_{\mathbf{n}^{\prime}}\right)=\delta\left(\theta_{\mathbf{n}}-\theta_{\mathbf{n}}^{\prime}\right)\delta\left(\varphi_{\mathbf{n}}-\varphi_{\mathbf{n}}^{\prime}\right)/\sin\theta_{\mathbf{n}}$. These commutation relations, the expression of the medium-assisted electric field $\hat{\mathbf{E}}^{(M)}_{\omega}$ and the expression of the scattering-assisted electric field $\hat{\mathbf{E}}^{(S)}_{\omega}$ guarantee the canonical commutation relations for the electromagnetic field and the bath oscillator fields describing the dielectric objects [15].
The bosonic field operators $\hat{\mathbf{f}}_{\omega}(\mathbf{r})$ and $\hat{g}_{\omega\mathbf{n}\nu}$ diagonalize the electromagnetic environment Hamiltonian,
$\hat{H}_{E}=\hat{H}^{(M)}_{E}+\hat{H}^{(S)}_{E}$ (12)
where
$\hat{H}^{(M)}_{E}=\int_{0}^{\infty}d\omega\hbar\omega\int_{V}d^{3}\,\mathbf{r}\,\hat{\mathbf{f}}_{\omega}^{\dagger}(\mathbf{r})\cdot\hat{\mathbf{f}}_{\omega}(\mathbf{r}),$ (13a)
$\hat{H}^{(S)}_{E}=\int_{0}^{\infty}d\omega\hbar\omega\,\oint do_{\mathbf{n}}\sum_{\nu}\hat{g}_{\omega\mathbf{n}\nu}^{\dagger}\hat{g}_{\omega\mathbf{n}\nu},$ (13b)
are, respectively, the contribution of the medium and scattering-assisted fields. The operators $\{\hat{\mathbf{f}}_{\omega}^{\dagger}$, $\hat{\mathbf{f}}_{\omega}\}$ and $\{\hat{g}_{\omega\mathbf{n}\nu}^{\dagger}$, $\hat{g}_{\omega\mathbf{n}\nu}\}$ are the creation and annihilation operators of polaritonic excitations in two independent bosonic reservoirs, the medium-assisted and scattering-assisted bosonic reservoirs, respectively.
The fundamental integral identity [15]
$\mathcal{A}_{\omega}(\mathbf{r},\mathbf{r}^{\prime})+\mathcal{B}_{\omega}(\mathbf{r},\mathbf{r}^{\prime})=\frac{\hbar\omega^{2}}{\pi\varepsilon_{0}c^{2}}\operatorname{Im}\left[\mathcal{G}_{\omega}\left(\mathbf{r},\mathbf{r}^{\prime}\right)\right]$ (14)
holds, where
$\mathcal{A}_{\omega}(\mathbf{r},\mathbf{r}^{\prime})$ $=\int_{V}d^{3}\mathbf{r^{\prime\prime}}\,\mathcal{G}_{e}(\mathbf{r},\mathbf{r^{\prime\prime};\omega})\cdot\mathcal{G}_{e}^{*T}\left(\mathbf{r}^{\prime},\mathbf{r^{\prime\prime}};\omega\right),$ (15a)
$\mathcal{B}_{\omega}(\mathbf{r},\mathbf{r}^{\prime})$ $=\oint do_{\mathbf{n}}\sum_{\nu}\mathbf{E}_{\omega\mathbf{n}\nu}(\mathbf{r})\mathbf{E}_{\omega\mathbf{n}\nu}^{*}\left(\mathbf{r}^{\prime}\right).$ (15b)
From this property it follows that
${\left[\hat{\mathbf{E}}_{\omega}(\mathbf{r}),\hat{\mathbf{E}}_{\omega^{\prime}}^{\dagger}\left(\mathbf{r}^{\prime}\right)\right]}=\frac{\hbar\omega^{2}}{\pi\varepsilon_{0}c^{2}}\operatorname{Im}\left[\mathcal{G}_{\omega}\left(\mathbf{r},\mathbf{r}^{\prime}\right)\right]\delta\left(\omega-\omega^{\prime}\right),$ (16)
which is the standard electric field commutator expressed in terms of the dyadic Green function of the dielectric.
III Bright modes of the electromagnetic environment and reduced Hamiltonian
We now represent the vector field operators $\{\hat{\mathbf{f}}_{\omega}(\mathbf{r})\}$ and the scalar field operators $\{\hat{g}_{\omega\mathbf{n}\nu}\}$ through the emitter-centered modes (e.g. [18], [19], [1]). Following Ref. [19], we aim at eliminating the degrees of freedom of the medium-assisted reservoir and the scattering-assisted reservoir that are not excited by the quantum emitters. Hence, starting from Eqs. 2, 3, and 4, we introduce the auxiliary interaction operator $\hat{F}_{j}^{(M)}$ describing the coupling between the $j$-th emitter and the the medium-assisted reservoir
$\hat{F}_{j}^{(M)}=\int_{0}^{\infty}d\omega\int_{V}d^{3}\mathbf{r}\,[\mathbb{u}_{j}\cdot\mathcal{G}_{e}(\mathbf{r}_{j},\mathbf{r},\omega)\cdot\hat{\mathbf{f}}_{\omega}(\mathbf{r}_{j})+h.c.].$ (17)
Similarly, from Eqs. 2, 3, and 10, we introduce the auxiliary interaction operator $\hat{F}_{j}^{(S)}$, describing the coupling between the $j$-th emitter and the scattering-assisted reservoir
$\hat{F}_{j}^{(S)}=\int_{0}^{\infty}d\omega\oint d{o}_{\mathbb{n}}\,[\sum_{\nu}\,\mathbb{u}_{j}\cdot\mathbf{E}_{\omega\mathbf{n}\nu}(\mathbf{r}_{j})\hat{g}_{\omega\mathbf{n}\nu}+h.c.]\,.$ (18)
The interaction Hamiltonian 2 can then be recast as
$\hat{H}_{I}=-\sum_{j=1}^{N}\hat{\mu}_{j}\hat{F}_{j}$ (19)
where
$\hat{F}_{j}=\hat{F}_{j}^{(M)}+\hat{F}_{j}^{(S)}.$ (20)
III.1 Medium-Assisted Reservoir
The structure of the expression 17 of $\hat{F}^{(M)}_{j}$ suggests introducing the monochromatic scalar bosonic operator
$\hat{A}_{j}(\omega)=\int_{V}d^{3}\mathbf{r}\,\bm{\alpha}_{j}(\mathbf{r};{\omega})\cdot\hat{\mathbf{f}}_{\omega}{(\mathbf{r})}$ (21)
with $j=1,N$, where
$\bm{\alpha}_{j}(\mathbf{r};{\omega})=\mathbb{u}_{j}\cdot\mathcal{G}_{e}(\mathbf{r}_{j},\mathbf{r};\omega){,}$ (22)
so that
$\hat{F}^{(M)}_{j}=\int_{0}^{\infty}d\omega[\hat{A}_{j}(\omega)+h.c.]\,.$ (23)
The operators $\{\hat{A}_{i}(\omega)\}$ obey
${\left[\hat{A}_{i}(\omega),\hat{A}_{j}^{\dagger}(\omega^{\prime})\right]}=M_{ij}(\omega)\delta\left(\omega-\omega^{\prime}\right)$ (24)
where
$M_{ij}(\omega)=\int_{V}d^{3}\mathbf{r}\,\bm{\alpha}_{i}(\mathbf{r};{\omega})\cdot\bm{\alpha}_{j}^{*}(\mathbf{r};{\omega}).$ (25)
The $N\times N$ matrix $M=[M_{ij}]$ is a complex Gram matrix associated with the inner product $\int_{V}d\mathbf{r}^{3}\mathbb{a}(\mathbf{r})\cdot\mathbb{b}^{*}(\mathbf{r})$. By construction $M$ is Hermitian. The vector fields $\{\bm{\alpha}_{j}(\mathbf{r};\omega)\}$ are linearly independent, and then $M$ is positive definite. Since these fields are not mutually orthogonal with respect to the inner product above, $M$ is, in general, dense. Using 22, we obtain
$M_{ij}(\omega)=\int_{V}{d^{3}\mathbf{r}\,\mathbb{u}_{i}\cdot\mathcal{G}_{e}(\mathbf{r}_{i},\mathbf{r}};\omega)\mathcal{G}_{e}^{*T}(\mathbf{r}_{j},\mathbf{r};\omega)\cdot\mathbb{u}_{j}.$ (26)
We now seek a representation of $\hat{H}_{E}^{(M)}$ in terms of the scalar operators $\{\hat{A}_{j}(\omega)\}$. The vector fields $\{\bm{\alpha}_{j}(\mathbf{r};\omega)\}$ are, in general, not orthogonal with respect to the above inner product; consequently, the basis $\{\hat{A}_{j}(\omega)\}$ does not diagonalize $\hat{H}_{E}^{(M)}$. To obtain a diagonal form, we generate the orthonormal set of vector fields $\{\bm{\chi}_{i}(\mathbf{r};\omega)\}$ by linear transformation
$\bm{\chi}_{i}(\mathbf{r};\omega)=\sum_{j=1}^{N}U_{ij}(\omega)\bm{\alpha}_{j}(\mathbf{r};\omega)$ (27)
where the $N\times N$ matrix $U=[U_{ij}]$ is such that $UMU^{\dagger}=I$ and II is the identity matrix. Completing $\{\bm{\chi}_{i}(\mathbf{r};\omega)\}$ with an orthonormal set $\{\bm{\chi}^{\mathrm{dark}}_{m}(\mathbf{r})\}$, we express the field operator $\hat{\mathbf{f}}_{\omega}(\mathbf{r})$ as
$\hat{\mathbf{f}}_{\omega}(\mathbf{r})=\sum_{j=1}^{N}\bm{\chi}_{j}^{*}(\mathbf{r};\omega)\hat{C}_{j}(\omega)+\sum_{m}[\bm{\chi}_{m}^{dark}(\mathbf{r})]^{*}\hat{C}_{m}^{dark}(\omega),$ (28)
where the new scalar bosonic operator $\hat{C}_{i}(\omega)$ is given by
$\hat{C}_{i}(\omega)=\sum_{j=1}^{N}U_{ij}(\omega)\hat{A}_{j}(\omega).$ (29)
Note that every vector field $\bm{\chi}_{m}^{dark}(\mathbf{r})$ does not couple to the emitters: the vector fields $\{\bm{\chi}_{i}(\mathbf{r};\omega)\}$ are the bright modes of the medium assisted field, and the vector fields $\{\bm{\chi}_{m}^{dark}(\mathbf{r})\}$ are the dark modes. The scalar bosonic operators $\hat{C}_{i}(\omega)$ are the medium-assisted bright bosonic operators of the system and $\hat{C}_{m}^{dark}(\omega)$ are the dark bosonic operators. The matrix $U$ can be chosen in various ways [19]. For our purposes, it is convenient to choose $U=M^{-1/2}$.
Monochromatic bosonic operators $\{\hat{C}_{i}(\omega)\}$ satisfy
${\left[\hat{C}_{i}(\omega),\hat{C}_{j}^{\dagger}(\omega^{\prime})\right]}=\delta_{ij}\delta\left(\omega-\omega^{\prime}\right).$ (30)
The bosonic operators $\hat{A}_{i}(\omega)$ are related to $\hat{C}_{i}(\omega)$ through the equations
$\hat{A}_{i}(\omega)=\sum_{j=1}^{N}G^{(M)}_{ij}(\omega)\hat{C}_{j}(\omega),$ (31)
where the $N\times N$ matrix $G^{(M)}=[G^{(M)}_{ij}$] is the inverse of the matrix $U$, $G^{(M)}={M}^{1/2}$. Using this relation, we express $\hat{F}^{(M)}_{i}$ through the bright bosonic operators $\{\hat{C}_{j}\}$,
$\hat{F}_{i}^{(M)}=\sum_{j=1}^{N}\int_{0}^{\infty}d\omega[G_{ij}^{(M)}(\omega)\hat{C}_{j}(\omega)+h.c.].$ (32)
Using 28, we obtain for $\hat{H}_{E}^{(M)}$
$\hat{H}_{E}^{(M)}=\sum_{j=1}^{N}\int_{0}^{\infty}d\omega\hbar\omega\hat{C}_{j}^{\dagger}(\omega)\hat{C}_{j}(\omega)\\ +\sum_{m}\int_{0}^{\infty}d\omega\hbar\omega\sum_{m}\hat{C}_{m}^{dark\dagger}({\omega})\hat{C}_{m}^{dark}({\omega}).$ (33)
III.2 Scattering-Assisted Reservoir
As in the medium-assisted case, the expression 18 of $\hat{F}_{j}^{(S)}$ suggests introducing the scalar monochromatic bosonic operator
$\hat{B}_{j}(\omega)=\oint do_{\mathbf{n}}\sum_{\nu}\,{\beta}_{j}(\mathbf{n},\nu;\omega)\hat{g}_{\omega\mathbf{n}\nu}$ (34)
with $j=1,N$, where
${\beta}_{j}(\mathbf{n},\nu;\omega)=\mathbb{u}_{j}\cdot\mathbf{E}_{\omega\mathbf{n}\nu}(\mathbf{r}_{j}),$ (35)
so that
$\hat{F}^{(S)}_{j}=\int_{0}^{\infty}d\omega[\hat{B}_{j}(\omega)+h.c.]\,.$ (36)
The operators $\{\hat{B}_{i}(\omega)\}$ obey
${\left[\hat{B}_{i}(\omega),\hat{B}_{j}^{\dagger}(\omega^{\prime})\right]}=S_{ij}(\omega)\delta\left(\omega-\omega^{\prime}\right)$ (37)
where
$S_{ij}(\omega)=\oint do_{\mathbf{n}}\,\mathbf{u}_{i}\cdot\sum_{\nu}\mathbf{E}_{\omega\mathbf{n}\nu}(\mathbf{r}_{i})\mathbf{E}_{\omega\mathbf{n}\nu}^{*}(\mathbf{r}_{j})\cdot\mathbb{u}_{j}.$ (38)
The $N\times N$ matrix $S=[S_{ij}]$ is also a complex Gram matrix with respect to the scalar product $\int do_{\mathbf{n}}\sum_{\nu}{f}(\mathbf{n},\nu){g}^{*}(\mathbf{n},\nu)$. Therefore, $S$ is Hermitian by construction; since the vectors $\{\beta_{j}(\mathbf{n},\nu,\omega)\}$ are linearly independent, it is also positive definite. Proceeding as in the medium-assisted case, we introduce the set of functions $\left\{\xi_{j}(\mathbb{n},\nu;\omega)\right\}$, orthonormal with respect to the above scalar product. They are given by
$\xi_{i}(\mathbb{n},\nu;\omega)=\sum_{j=1}^{N}V_{ij}{\beta}_{j}(\mathbf{n},\nu;\omega)$ (39)
where the $N\times N$ matrix $V=[V_{ij}]$ is such that $VSV^{\dagger}=I$. Regarding the bright modes of the field assisted by the medium, we choose $V=S^{-1/2}$. Completing the set $\{{\xi}_{i}(\mathbb{n},\nu;\omega)\}$ with an orthonormal set $\{{\xi}^{\mathrm{dark}}_{m}\}$, we express the field operator $\hat{g}_{\omega\mathbf{n}\nu}$ as
$\hat{g}_{\omega\mathbf{n}\nu}=\sum_{j=1}^{N}\xi_{j}^{*}(\mathbb{n},\nu,\omega)\hat{D}_{j}(\omega)+\sum_{m}[{\xi}_{m}^{dark}]^{*}\hat{D}_{m}^{dark}(\omega),$ (40)
where the bosonic operators $\{\hat{D}_{i}(\omega)\}$ are given by
$\hat{D}_{i}(\omega)=\sum_{j=1}^{N}V_{ij}(\omega)\hat{B}_{j}(\omega).$ (41)
As for the vector fields $\{\bm{\chi}_{m}^{dark}\}$, the scalar fields $\{\xi_{m}^{dark}\}$ do not pair up with the emitters; $\{\xi_{i}\}$ are the bright modes of the scattering-assisted field and $\{\hat{D}_{i}\}$ are the corresponding bright operators. The scalar bosonic operators $\{\hat{B}_{i}(\omega)\}$ are related to $\{\hat{D}_{i}(\omega)\}$ through the equations
$\hat{B}_{i}(\omega)=\sum_{j=1}^{N}G_{ij}^{(S)}(\omega)\hat{D}_{j}(\omega),$ (42)
where the $N\times N$ matrix $G^{(S)}=[G^{(S)}_{ij}$] is the inverse of the matrix $V$, $G^{(S)}=S^{1/2}$. Using this relation, we express $\hat{F}_{i}^{(S)}$ through the bright bosonic operators $\{\hat{D}_{j}(\omega)\}$,
$\hat{F}_{i}^{(S)}=\sum_{j=1}^{N}\int_{0}^{\infty}d\omega[G_{ij}^{(S)}(\omega)\hat{D}_{j}(\omega)+h.c.].$ (43)
Using 39, we obtain for $\hat{H}_{E}^{(S)}$
$\hat{H}_{E}^{(S)}=\sum_{j=1}^{N}\int_{0}^{\infty}d\omega\hbar\omega\hat{D}_{j}^{\dagger}(\omega)\hat{D}_{j}(\omega)\\ +\sum_{m}\int_{0}^{\infty}d\omega\hbar\omega\sum_{m}\hat{D}_{m}^{dark\dagger}({\omega})\hat{D}_{m}^{dark}({\omega}).$ (44)
III.3 Reduced Hamiltonian
The medium-assisted and scattering-assisted dark modes are decoupled from the rest of the system, they do not affect dynamics of the quantum emitters and can be dropped from the expressions of $\hat{H}_{E}^{(M)}$ and $\hat{H}_{E}^{(S)}$. The overall system, quantum emitters + bright modes of the electromagnetic environment, is described by the reduced Hamiltonian
$\hat{H}_{red}=\hat{H}_{A}+\hat{H}_{E}^{bright}+\hat{H}_{I}$ (45)
where
$\hat{H}_{E}^{bright}=\sum_{i=1}^{N}\int_{0}^{\infty}d\omega\hbar\omega\left[\hat{C}_{i}^{\dagger}(\omega)\hat{C}_{i}(\omega)+\hat{D}_{i}^{\dagger}(\omega)\hat{D}_{i}(\omega)\right]$ (46)
and
$\hat{H}_{I}=\sum_{i,j=1}^{N}\hat{\mu}_{i}\left[\int_{0}^{\infty}d\omega G_{ij}^{(M)}(\omega)\hat{C}_{j}(\omega)+h.c.\right]+\\ \sum_{i,j=1}^{N}\hat{\mu}_{i}\left[\int_{0}^{\infty}d\omega G_{ij}^{(S)}(\omega)\hat{D}_{j}(\omega)+h.c.\right].$ (47)
In summary, the set of quantum emitters behaves as an open quantum system coupled to two independent reservoirs: the medium-assisted reservoir and the scattering-assisted reservoir. Each reservoir consists of $N$ independent set of bosonic modes. The medium-assisted reservoir is characterized by the coupling matrix ${G}^{(M)}(\omega)=M(\omega)^{1/2}$, and the scattering-assisted reservoir by ${G}^{(M)}(\omega)=S(\omega)^{1/2}$. In Appendix A, we provide the expression of the electric field operator due to the bright modes.
IV Spectral density matrices
The matrices $G^{(M)}(\omega)$ and $G^{(S)}(\omega)$ fully characterize the interaction of the quantum emitters with the electromagnetic environment, including their mutual interactions. It is convenient to recast these coupling matrices in terms of two matrices that generalize the concepts of medium-assisted spectral density and scattering-assisted spectral density, as originally introduced in [1] and [2], for a single quantum emitter, to configurations of multiple quantum emitters.
Let us introduce the $N\times N$ diagonal matrix $P$ of the transition dipole moments of the quantum emitters as
$P=\text{diag}(\mu_{1},\mu_{2},...,\mu_{N})$ (48)
where $\mu_{j}$ is the transition dipole moment of the $j$-th quantum emitter. The medium-assisted spectral density matrix $\mathcal{J}^{(M)}$ and the scattering-assisted spectral density matrix $\mathcal{J}^{(S)}$ are defined as
$\mathcal{J}^{(M)}(\omega)=\frac{1}{\hbar^{2}}P\,[{G}^{(M)}(\omega)]^{2}\,P,$ (49a)
$\mathcal{J}^{(S)}(\omega)=\frac{1}{\hbar^{2}}P\,[{G}^{(S)}(\omega)]^{2}\,P,$ (49b)
so that, equivalently,
${G}^{(M)}(\omega)$ $=\hbar[P^{-1}\mathcal{J}^{(M)}(\omega)P^{-1}]^{1/2},$ (50a)
${G}^{(S)}(\omega)$ $=\hbar[P^{-1}\mathcal{J}^{(S)}(\omega)P^{-1}]^{1/2}.$ (50b)
Since $G^{(M)}=M^{1/2}$ and $G^{(S)}=S^{1/2}$, we obtain
$\mathrm{\mathcal{J}}_{ij}^{(M)}(\omega)$ $=\frac{\mu_{i}\mu_{j}}{\hbar^{2}}M_{ij},$ (51a)
$\mathrm{\mathcal{J}}_{ij}^{(S)}(\omega)$ $=\frac{\mu_{i}\mu_{j}}{\hbar^{2}}S_{ij},$ (51b)
and using 26 and 38,
$\mathrm{\mathcal{J}}_{ij}^{(M)}(\omega)$ $=\frac{\mu_{i}\mu_{j}}{\hbar^{2}}\int_{V}{d^{3}\mathbf{r}\,\mathbb{u}_{i}\cdot\mathcal{G}_{eq}(\mathbf{r}_{i},\mathbf{r}})\mathcal{G}_{eq}^{*T}(\mathbf{r}_{j},\mathbf{r})\cdot\mathbb{u}_{j},$ (52a)
$\mathrm{\mathcal{J}}_{ij}^{(S)}(\omega)$ $=\frac{\mu_{i}\mu_{j}}{\hbar^{2}}\oint do_{\mathbf{n}}\,\sum_{\nu}\mathbf{u}_{i}\cdot\mathbf{E}_{\omega\mathbf{n}\nu}(\mathbf{r}_{i})\mathbf{E}_{\omega\mathbf{n}\nu}^{*}(\mathbf{r}_{j})\cdot\mathbb{u}_{j}.$ (52b)
Both spectral density matrices are complex, Hermitian, and positive definite. The elements $\mathrm{\mathcal{J}}_{ij}^{(M)}(\omega)$ and $\mathrm{\mathcal{J}}_{ij}^{(S)}(\omega)$ are not independent; in fact, as a consequence of 14
$\mathrm{\mathcal{J}}_{ij}^{(M)}(\omega)+\mathrm{\mathcal{J}}_{ij}^{(S)}(\omega)=\frac{1}{2\pi}\Gamma_{ij}(\omega)$ (53)
where
$\Gamma_{ij}(\omega)=\frac{2\omega^{2}}{\hbar\varepsilon_{0}c^{2}}(\mu_{i}\mathbb{u}_{i})\cdot\operatorname{Im}\left[\mathcal{G}_{\omega}\left(\mathbf{r}_{i}.\mathbf{r}_{j}\right)\right]\cdot(\mu_{j}\mathbb{u}_{j}).$ (54)
The matrix $\Gamma=[\Gamma_{ij}]$ is real, symmetric, and positive definite.
The medium-assisted spectral density and the scattering-assisted spectral density have a direct interpretation within classical electrodynamics. Consider the classical electromagnetic field generated, in the presence of dielectric objects, by $N$ electric dipoles with dipole moments $\mu_{j}\mathbf{u}_{j}$, located at ${\mathbf{r}}_{j}$ with $j=1,N$ and oscillating at frequency $\omega$. We find that $\frac{\pi}{2}\hbar\omega\sum_{i,j=1}^{N}\mathcal{J}^{(M)}_{ij}(\omega)$ is equal to the average electromagnetic power absorbed by the dielectric bodies; $\frac{1}{4}(\hbar\omega)\sum_{i,j=1}^{N}\Gamma_{ij}(\omega)$ is equal to the averaged electromagnetic power emitted by the set of electric dipoles. Then, $\frac{\pi}{2}\hbar\omega\sum_{i,j=1}^{N}\mathcal{J}^{(S)}_{ij}(\omega)$ is equal to the average electromagnetic power radiated toward infinity. Accordingly, Eq. 53 represents the statement of the Poynting theorem for the investigated scenario. See Appendix C for a full discussion.
V Correlator matrix and surrogate environment
The dynamic of $N$ quantum emitters coupled to $N$ medium-assisted bosonic modes characterized by the spectral density matrix $\mathcal{J}^{(M)}(\omega)$ and to $N$ scattering-assisted bosonic modes characterized by the spectral density matrix $\mathcal{J}^{(S)}(\omega)$ can, in principle, be computed numerically with a variety of methods [20, 21]. However, this is often computationally demanding. Under specific conditions, analogous to the single quantum emitter [1], [2] case, the actions of the $2N$ bosonic modes can be represented by $N$ bosonic modes characterized by an equivalent $N\times N$ spectral density matrix, in the same spirit of [22, 23].
We assume that: (i) the initial quantum state of the entire system is a product state, that is, $\hat{\rho}(0)=\hat{\rho}_{A}(0)\otimes\hat{\rho}_{E}(0)$, where $\hat{\rho}(0)$, $\hat{\rho}_{A}(0)$ and $\hat{\rho}_{E}(0)$ are the density operators of the entire system, of the set of quantum emitters and of the electromagnetic environment, respectively; (ii) the initial state of the electromagnetic environment is Gaussian. Then, the evolution of the reduced density operator of the set of quantum emitters $\hat{\rho}_{A}(t)=\text{Tr}_{E}\left[\rho(t)\right]$ depends only on the expectation values ${F}_{j}(t)$ and the two-time correlation functions ${C}_{ij}(t+\tau;t)$ (e.g. [24]) of the environment interaction operators $\hat{F}_{j}$ with i,j=1,2,,Ni,j=1,2,...,N,
${F}_{j}(t)$ $=\text{Tr}_{E}\left[\hat{F}_{j}(t)\hat{\rho}_{E}(0)\right],$ (55a)
${C}_{ij}(t+\tau;t)$ $=\text{Tr}_{E}\left[{\hat{F}}_{i}(t+\tau){\hat{F}}_{j}(t)\hat{\rho}_{E}(0)\right],$ (55b)
where
$\hat{{F}}_{j}(t)=\hat{U}^{\dagger}_{E}(t)\hat{F}_{j}\hat{U}_{E}(t),$ (56)
and $\hat{U}_{E}(t)=\exp(-i\hat{H}_{E}^{bright}t/\hbar)$ is the free evolution operator of the bright bosonic modes of the electromagnetic environment. Using this property, it is possible to introduce an equivalent surrogate environment with only $N$ bosonic modes that reproduces the same reduced dynamics of the quantum emitters.
The expectation values and the two-time correlation functions of the interaction operators are given by:
${F}_{j}(t)={F}^{(M)}_{j}(t)+{F}^{(S)}_{j}(t)$ (57)
and
$C_{ij}(t+\tau;t)=C^{(M)}_{ij}(t+\tau;t)+C^{(S)}_{ij}(t+\tau;t)+\\ {F}^{(M)}_{i}(t+\tau){F}^{(S)}_{j}(t)+{F}^{(S)}_{i}(t+\tau){F}^{(M)}_{j}(t),$ (58)
where ${F}^{(\alpha)}_{i}(t)=\text{Tr}_{E}\left[\hat{F}^{(\alpha)}_{i}(t)\hat{\rho}_{E}(0)\right]$ and ${C}^{(\alpha)}_{ij}(t+\tau;t)=\text{Tr}_{E}\left[{\hat{F}}^{(\alpha)}_{i}(t+\tau){\hat{F}}^{(\alpha)}_{j}(t)\hat{\rho}_{E}(0)\right]$ with $\alpha=M,S$. If these expectation values vanish, we have
$C_{ij}(t)=C^{(M)}_{ij}(t)+C^{(S)}_{ij}(t).$ (59)
This is the case, for instance, when both reservoirs are initially in vacuum, thermal, or number states. In such situations, the influence of the electromagnetic environment on the emitter ensemble is fully characterized by the $N\times N$ correlator matrix $C(t)=[C_{ij}(t)]$.
V.1 Correlators for Thermal Initial States
We now consider scenarios in which: (a) the initial quantum state of the electromagnetic environment is a product state, that is, $\hat{\rho}_{E}(0)=\hat{\rho}_{E}^{(M)}(0)\otimes\hat{\rho}_{E}^{(S)}$, where $\hat{\rho}_{E}^{(M)}(0)$ and $\hat{\rho}_{E}^{(S)}(0)$ are the initial density operators of the medium-assisted reservoir and of the scattering-assisted reservoir, respectively; (b) the two reservoirs are initially in thermal quantum states with temperatures $T_{0}^{(M)}$ and $T_{0}^{(S)}$. Under these assumptions, the correlation function $C_{ij}^{(\alpha)}(t)$ with $\alpha=M,S$ becomes (see Appendix C for details):
$C_{ij}^{(\alpha)}(t)=\frac{\hbar^{2}}{\mu_{i}\mu_{j}}\int_{0}^{\infty}\mathrm{d}\omega[(1+n_{\omega}^{(\alpha)})\mathcal{J}_{ij}^{(\alpha)}(\omega)e^{-i\omega t}+\\ n_{\omega}^{(\alpha)}\mathcal{J}_{ij}^{(\alpha)*}(\omega)e^{+i\omega t}]$ (60)
where
$n_{\omega}^{(\alpha)}=\frac{1}{e^{\beta_{\alpha}\hbar\omega}-1},$ (61)
and $\beta_{\alpha}=1/(k_{B}T_{0}^{(\alpha)})$. It is convenient to rewrite 60 as
$C^{(\alpha)}_{ij}(t)=\frac{\hbar^{2}}{\mu_{i}\mu_{j}}\int_{-\infty}^{+\infty}\mathrm{d}\omega\,\mathrm{J}^{(\alpha)}_{ij}(\omega;\beta_{\alpha})e^{-i\omega t},$ (62)
where $\mathrm{J}^{(\alpha)}_{ij}(\omega;\beta_{\alpha})$, defined for $\infty<\omega<+\infty$, is
$\quad\;\mathrm{J}^{(\alpha)}_{ij}(\omega;\beta_{\alpha})=\begin{cases}(1+n_{\omega}^{(\alpha)})\mathcal{J}_{ij}^{(\alpha)}(\omega)&\text{for}\,\omega\geq 0,\\ n_{|\omega|}^{(\alpha)}\mathcal{J}_{ij}^{(\alpha)*}(|\omega|)&\text{for}\,\omega\leq 0.\end{cases}$ (63)
The function $\mathrm{J}^{(\alpha)}_{ij}(\omega;\beta_{\alpha})$ is continuous at $\omega=0$ because the spectral density matrices go to zero for $\omega\rightarrow 0$ at least as $\omega^{2}$. Eventually, the correlation function $C_{ij}(t)$ can be expressed as
$C_{ij}(t)=\frac{\hbar^{2}}{\mu_{i}\mu_{j}}\int_{-\infty}^{+\infty}d\omega\mathrm{J}^{eff}_{ij}(\omega;\beta_{M},\beta_{S})e^{-i\omega t},$ (64)
where
$\mathrm{J}^{\text{eff}}_{ij}(\omega;\beta_{M},\beta_{S})=\mathrm{J}^{(M)}_{ij}(\omega;\beta_{M})+\mathrm{J}^{(S)}_{ij}(\omega;\beta_{S}).$ (65)
When the medium-assisted reservoir and the scattering-assisted reservoir are at the same temperature $T_{0}$ we obtain
$\quad\;\mathrm{J}^{\text{eff}}_{ij}(\omega;\beta)=\begin{cases}\frac{1}{2\pi}(1+n_{\omega})\Gamma_{ij}(\omega)&\text{for}\,\omega\geq 0,\\ \frac{1}{2\pi}n_{|\omega|}\Gamma_{ij}^{*}(|\omega|)&\text{for}\,\omega\leq 0,\end{cases}$ (66)
where now
$n_{\omega}=\frac{1}{e^{\beta\hbar\omega}-1},$ (67)
and $\beta=1/(k_{B}T_{0})$. In the zero-temperature limit, expression 66 reduces to
$\quad\;\mathrm{J}^{\text{eff}}_{ij}(\omega;\infty)=\begin{cases}\frac{1}{2\pi}\Gamma_{ij}(\omega)&\text{for}\,\omega\geq 0,\\ 0&\text{for}\,\omega\leq 0.\end{cases}$ (68)
V.2 Surrogate Environment Initially in Vacuum Quantum State
Let $\mathrm{J}^{(M)}(\omega;\beta_{M})$ and $\mathrm{J}^{(S)}(\omega;\beta_{S})$ denote the $N\times N$ complex matrices with elements $\mathrm{J}^{(M)}_{ij}(\omega;\beta_{M})$ and $\mathrm{J}^{(S)}_{ij}(\omega;\beta_{S})$, respectively. We call these matrices temperature-dependent spectral density matrices of the medium and scattering-assisted reservoirs, respectively. These matrices are defined for $\infty<\omega<+\infty$, Hermitian, and positive definite. We now introduce the effective spectral density matrix
$\mathrm{J}^{\text{eff}}(\omega;\beta_{M},\beta_{S})=\mathrm{J}^{(M)}(\omega;\beta_{M})+\mathrm{J}^{(S)}(\omega;\beta_{S}),$ (69)
which characterizes the overall electromagnetic environment. This matrix is also defined for $\infty<\omega<+\infty$, complex, Hermitian, and positive definite.
By the equivalence established above, the reduced dynamics of the emitters can be computed by considering the surrogate Hamiltonian
$\hat{H}^{sur}=\hat{H}_{A}+\hat{H}_{E}^{sur}+\hat{H}_{I}^{sur}$ (70)
where $\hat{H}_{E}^{sur}$ is the bare Hamiltonian of a surrogate environment consisting of $N$ bosonic reservoirs, with positive and negative frequencies, initially in the vacuum state,
$\hat{H}_{E}^{sur}=\sum_{j=1}^{N}\int_{-\infty}^{+\infty}d\omega\hbar\omega\hat{a}_{j}^{\dagger}(\omega)\hat{a}_{j}(\omega);$ (71)
here $\hat{a}_{j}(\omega)$ and $\hat{a}_{j}^{\dagger}(\omega)$, with $j=1,N$, are the annihilation and creation operators of the bosonic modes of the surrogate environment. The interaction Hamiltonian $\hat{H}_{I}^{sur}$ between the quantum emitters and the surrogate environment is given by
$\hat{H}_{I}^{surr}=-\sum_{i,j=1}^{N}\hat{\mu}_{i}\int_{-\infty}^{+\infty}d\omega[G^{\text{eff}}_{ij}(\omega)\hat{a}_{j}(\omega)+h.c.],$ (72)
where $G^{\text{eff}}_{ij}(\omega)$ is the $ij$-th element of the $N\times N$ coupling matrix $G^{\text{eff}}(\omega)$ given by
${G}^{\text{eff}}(\omega)=\hbar[D^{-1}{J}^{\text{eff}}(\omega;\beta_{M},\beta_{M})D^{-1}]^{1/2}.$ (73)
When the medium-assisted reservoir and the scattering-assisted reservoir are initially in the vacuum state, the effective spectral density $\mathrm{J}_{ij}^{\text{eff}}(\omega)$ is given by 68. In the literature relevant to multiple quantum emitters that interact with dielectric objects based on the Langevin noise formalism, this expression is used [8]. However, as already pointed out in [1], [2] for a single quantum emitter, the conditions under which 68 remains valid are not always clearly stated. In fact, when the two reservoirs are initially in thermal quantum states, the expression 68 is not more valid. If the temperatures of the two reservoirs are the same, the elements of the effective spectral density matrix are given by 66. When the initial temperatures are different, the elements of the effective spectral density matrix are given by 65.
Figure 1: Two equal quantum emitters with dipole moment $\mu$ interact with a Drude sphere with radius $a$, plasma frequency $\omega_{p}$ and damping rate $\nu$. Normalized spectral densities $\mathcal{J}^{(M)}_{ij}/\mathcal{J}_{0}$ and $\mathcal{J}^{(S)}_{ij}/\mathcal{J}_{0}$, for $i,j=1,2$, as a function of the normalized frequency $\omega/\omega_{p}$ with $k_{p}\,a=1$, $\nu/\omega_{p}=0.01$ where $k_{p}=\omega_{p}/c$. The two emitters are positioned as in the inset, at the same distance $d$ from the sphere center, $d/a=1.75$. The characteristic spectral density $\mathcal{J}_{0}$ is given by $\mathcal{J}_{0}=\frac{1}{6\pi^{2}}\frac{k_{p}}{\hbar\,\varepsilon_{0}}\left({\mu k_{p}}\right)^{2}$. The spectral density matrices are Hermitian.
Figure 2: Two equal quantum emitters with dipole moment $\mu$ interact interact with a nanostructure composed of two rods and a disk. The first rod has length $3a$, width $2a$, and height $a$; the second rod has length $4a$, width $2a$, and height $a$; the disk has radius $a$ and height $a$. All three particles are made of a Drude material with plasma frequency $\omega_{p}$ and damping rate $\nu$ such as $\nu/\omega_{p}=0.01$. The two quantum emitters are placed at the midpoint of the gaps between adjacent particles. Normalized spectral densities $\mathcal{J}^{(S)}_{ij}/\mathcal{J}_{0}$ and $\mathcal{J}^{(M)}_{ij}/\mathcal{J}_{0}$ versus $\omega/\omega_{p}$ for $i,j=1,2$ and $k_{p}\,a=1$ where $k_{p}=\omega_{p}/c$. The characteristic spectral density $\mathcal{J}_{0}$ is given by $\mathcal{J}_{0}=\frac{1}{6\pi^{2}}\frac{k_{p}}{\hbar\,\varepsilon_{0}}\left({\mu k_{p}}\right)^{2}$. The spectral density matrices are Hermitian.
VI Application to a system with two quantum emitters
In order to showcase the effects of an electromagnetic environment consisting of dispersive dielectric particles and vacuum on a system made of multiple quantum emitters, we examine the minimal setting of two quantum emitters, labeled 1 and 2. We consider two representative electromagnetic environments: a spherical particle in vacuum, as in [2], and a nanostructure composed of two rods and a disk in vacuum (see the insets of Figs. 1 and 2, respectively). The two quantum emitters are placed at the positions indicated by the red dots in the insets. All particles are homogeneous and their dielectric permittivity is described by the Drude model: $\varepsilon_{\omega}=[1-\omega_{p}^{2}/(\omega^{2}+i\nu\omega)]$ with plasma frequency $\omega_{p}$ and relaxation frequency $\nu$.
Within the modified Langevin noise formalism and the emitter-centered description, the electromagnetic environment is modeled by two continuous bosonic baths, medium (M)- and scattering (S)- assisted reservoirs, characterized by the spectral density matrices $\mathcal{J}^{(M)}(\omega)$ and $\mathcal{J}^{(S)}(\omega)$. The two reservoirs are initially in thermal quantum states at inverse temperatures $\beta_{\text{M}}$ and $\beta_{\text{S}}$. The two emitters interact through their dipole moments with the continuous sets of bosonic excitations of the electromagnetic environment.
Each emitter is modeled as a two-level system with transition frequency $\Omega_{i}$ and dipole moment magnitude $\mu_{i}$ ($i=1,2$). Hence, the bare quantum emitter Hamiltonian $\hat{H}_{A}$ is given by
$\hat{H}_{A}=\frac{\hbar\Omega_{1}}{2}\sigma_{z}^{(1)}+\frac{\hbar\Omega_{2}}{2}\sigma_{z}^{(2)},$ (74)
and the transition dipole moment operators are $\hat{\mu}_{1}=\mu_{1}\hat{\sigma}^{(1)}_{x}$, $\hat{\mu}_{2}=\mu_{2}\hat{\sigma}^{(2)}_{x}$; here $\sigma_{z}^{(i)}=|e^{(i)}\rangle\langle e^{(i)}|-|g^{(i)}\rangle\langle g^{(i)}|$ and $\sigma_{x}^{(i)}=|g^{(i)}\rangle\langle e^{(i)}|+|e^{(i)}\rangle\langle g^{(i)}|$ where $|g^{(i)}\rangle$ and $|e^{(i)}\rangle$ are the ground and excited states of the $i$-th quantum emitter. In the numerical examples below, we considered equal dipole moments $\mu_{1}=\mu_{2}\equiv\mu$.
As shown in Sec. V, leveraging the temperature-dependent spectral densities $\mathrm{J}^{(M)}(\omega;\beta_{M})$ and $\mathrm{J}^{(S)}(\omega;\beta_{S})$, the degrees of freedom of the medium-assisted and scattering-assisted baths are reduced to those of two distinct sets of surrogate continuous bosonic reservoirs initially in the vacuum quantum state and with an effective spectral density matrix $\mathrm{J}^{\text{eff}}(\omega;\beta_{M},\beta_{S})$. Furthermore, each emitter is coupled to both continuous reservoirs through direct coupling $G^{\text{eff}}_{ii}(\omega)$ and cross-coupling amplitudes $G^{\text{eff}}_{ij}(\omega)$. The coupling of two emitters with a common reservoir has been known to generate quantum correlated states of the composite system [25], similar to what has been studied in the field of cQED [26]. However, any real electromagnetic environment with linear constitutive relations is characterized by a number of reservoirs equal to the number of emitters: each emitter is simultaneously coupled to all of them. We study the non-equilibrium dynamics of the emitters for a set of physically relevant initial quantum states, including pure product states and maximally entangled Bell states.
We investigate the reduced dynamics of the two quantum emitters using the Matrix Product State approach, similarly to what is done with the TEDOPA method [23], to simulate the time evolution of the whole emitter + surrogate environment state governed by the effective Hamiltonian defined by Eqs. 70-74. Such methods have been shown to reproduce exact results and to provide corrections to more recent numerical techniques based on collisional models [27].
VI.1 Spectral Densities
We begin by analyzing the behavior of the normalized spectral densities $\mathcal{J}_{ij}^{(M)}(\omega)/\mathcal{J}_{0}$ and $\mathcal{J}_{ij}^{(S)}(\omega)/\mathcal{J}_{0}$ as a function of $\omega/\omega_{p}$ for $i,j\in\{1,2\}$, where
$\mathcal{J}_{0}=\frac{1}{6\pi^{2}}\frac{k_{p}}{\hbar\,\varepsilon_{0}}\left({\mu k_{p}}\right)^{2}$ (75)
and $k_{p}=\omega_{p}/c$. By definition, the characteristic spectral density $\mathcal{J}_{0}$ is equal to the value of the vacuum spectral density at the wavenumber $k_{p}$, and indicates the coupling strength between the quantum emitter and the vacuum field. Since the spectral density matrices are Hermitian, we omit the plots of $\mathcal{J}_{21}^{(M)}(\omega)$ and $\mathcal{J}_{21}^{(S)}(\omega)$ . We evaluated numerically the spectral density matrices by using a surface integral equation formulation, as described in Appendix D.
Figure 1 shows $\mathcal{J}^{(S)}_{ij}/\mathcal{J}_{0}$ and $\mathcal{J}^{(M)}_{ij}/\mathcal{J}_{0}$ for a Drude sphere with radius $a$ as a function of the normalized frequency $\omega/\omega_{p}$, assuming size parameter $k_{p}\,a=1$ and normalized damping rate $\nu/\omega_{p}=0.01$. The two emitters are positioned as in the inset, each at a distance $d$ from the center of the sphere, with $d/a=1.75$. In the low–frequency region around the dipolar plasmon resonance ($\omega/\omega_{p}\simeq 0.54$), the amplitude of the elements $\mathcal{J}^{(S)}_{ij}$ is greater than or equal to $\mathcal{J}^{(M)}_{ij}$. At higher frequencies, near higher-order plasmonic resonances, the amplitude of $\mathcal{J}^{(M)}_{ij}$ becomes significantly larger than that of $\mathcal{J}^{(S)}_{ij}$. For a comprehensive analysis of the role of the scattering modes in shaping the overall spectral density, see Ref. [2]. The bandwidths of the plasmonic modes are limited by a combination of radiative and material losses, as detailed in [2].
Figure 2 shows the corresponding normalized spectral densities for a nanostructure composed of two rods and a disk. The first rod has length $3a$, width $2a$, and height $a$; the second rod has length $4a$, width $2a$, and height $a$; the disk has radius $a$ and height $a$. The emitters are placed at the midpoint of the gaps between adjacent particles. All materials follow the same Drude model with $k_{p}\,a=1$ and a damping rate $\nu/\omega_{p}=0.01$. Also in this case the scattering-assisted spectral densities are typically greater than medium-assisted at low frequencies, while at low frequencies the material-assisted contribution dominates and is typically characterized by sharper peaks associated to higher multipolar scattering orders.
The effective spectral density matrix $\mathrm{J}^{\text{eff}}(\omega;\beta_{M},\beta_{S})$ completely characterizes the interaction of the two emitters with the surrogate bosonic environment through Eq. 73: it is the sum of temperature-dependent medium-assisted $\mathrm{J}^{(M)}(\omega;\beta_{M})$ and temperature-dependent scattering-assisted $\mathrm{J}^{(S)}(\omega;\beta_{S})$ spectral density matrices. They are proportional to the medium- and scattering-assisted spectral density matrices and depend on the temperatures through the Bose-Einstein distribution according to Eq. 66 .
VI.2 Entanglement Decay and Revivals
Since the early days of quantum information theory [28], quantum entanglement [29, 30] has been recognized as a key resource for a variety of tasks [31]. During the past two decades, its generation [25, 32, 33, 34, 35], sudden death [36, 37, 38, 39], birth [40], and degradation [41, 42, 43, 44] under multiple external environments [45] have been extensively investigated. In particular, if the quantum bipartite system undergoes non-Markovian dynamics [46, 47, 48, 49, 33, 35], it has been shown that the quantum state experiences revivals in its entanglement features. More recently, the generation of entangled states of pairs of quantum emitters using plasmonic [50], as well as photonic 1D waveguides [5] through bound states in the continuum [51, 52], has been studied.
[…]
Allowing the medium- and scattering-assisted reservoirs to be in thermal states at different temperatures, with a the material-assisted reservoir playing the role of the hot bath, see Fig. 4, leads to increased populations of the excited states of the emitters. This behavior is justified by the increased probability of photon absorption from the warmer reservoir, which partially counteracts emission.
The properties of the reduced states shown in Figs. 3 and 4 stem from the evolution of the initially pure entangled state of the composite system under its coupling to the photon reservoirs. As the system relaxes, the state becomes progressively mixed, while initial quantum correlations, such as entanglement, are progressively degraded. This process is illustrated in Fig. 5, where the entanglement negativity $\mathcal{N}(\hat{\rho}_{12})$ is plotted as a function of time for different temperatures of the thermal states of the electromagnetic environment and frequency detuning of the emitter transition frequencies. The detrimental impact of finite-temperature bath states on negativity is due to the combined effect of increased populations of excited emitter states and of the enhanced decay of two-body correlations, sustained by off-diagonal terms of $\hat{\rho}_{12}$. The Rabi-like interaction with the electromagnetic environment thus leads the composite state to a separable state $\mathcal{N}(\hat{\rho}_{12})=0$. This relaxation is governed by the interplay between direct- and cross- correlations, both of which extend over a broad frequency range due to the presence of the scattering-assisted reservoir.
At stronger coupling, the physics becomes richer. Figure 6 shows the behavior of $\mathcal{N}(\rho_{12})$ for different values of the dipole momentum $\mu$, inverse temperatures of the electromagnetic environment, and fixed detuning of the emitter frequencies. Negativity undergoes sudden death at short times, followed by revivals of decreasing amplitudes at later times, hallmark of non-Markovian effects [35], arising from emitter-photon correlations, which are present even in the RWA limit [46]. However, in the case of Eq. (72), Hamiltonian terms beyond RWA can be proved to reduce the amplitudes of revivals. In addition, the finite-temperature states of the medium-assisted reservoir noticeably curb the entanglement revivals.
The analysis of negativity is repeated in Fig. 7 for the configuration of Fig. 2, where two emitters interacts with two rods and a nano-disk, under equal, low temperatures of the medium and scattering-assisted reservoirs. Transient, short-time entanglement revivals exhibit a marked dependence on detuning: for zero detuning and moderate coupling, the revivals are enhanced at short times. Altogether, these results highlight how the detailed structure of the medium- and scattering-assisted spectral density matrices affects the degradation and revival of entanglement, suggesting concrete routes to optimize quantum correlations in structured nanophotonic environments.
Figure 3: Spherical particle, see Fig. 1. Relaxation dynamics of populations of quantum emitters: $p_{a}^{e}(t)$ is the population of the quantum emitter $a$, for $a=1,2$, in the excited state $|e\rangle$. The initial composite state has been chosen as $|\hat{\rho}_{12}(0)=\outerproduct{\Psi^{-}}{\Psi^{-}}$, while $\mu=10^{-3}\omega_{p}$ and $\beta_{S}=\beta_{M}=1000$. solid blue (red) curves denote the case $\Omega_{1}=0.54\omega_{p},\Omega_{2}=\omega_{p}$, while cyan (orange) are computed for $\Omega_{1}=\Omega_{2}=0.54\omega_{p}$.
Figure 4: Spherical particle, see Fig. 1. Relaxation dynamics of populations of the quantum emitters: $p_{a}^{e}(t)$ is the population of the quantum emitter $a$, for $a=1,2$, in the excited state $|e\rangle$. The initial composite state has been chosen as $|\rho_{12}(0)=\outerproduct{\Psi^{-}}{\Psi^{-}}$, while $\mu=10^{-3}\omega_{p}$ and $\Omega_{1}=\Omega_{2}=0.54\omega_{p}$. Solid blue (red) curves are computed for $\beta_{S}=\beta_{M}=1000$, and dashed curves denote the case $\beta_{M}=1,\beta_{S}=1000$.
Figure 5: Spherical particle, see Fig. 1. Time evolution of negativity $\mathcal{N}(\hat{\rho}_{12})$ with $\mu=10^{-3}\omega_{p}$, corresponding to the same initial state as in Figs. 3 and 4. Solid blue(red) lines denote the cases $\Omega_{1}=0.54\omega_{p},\Omega_{2}=\omega_{p}$ and $\Omega_{1}=\Omega_{2}=0.54\omega_{p}$ with equal inverse temperatures $\beta_{S}=\beta_{M}=1000$, respectively, while dashed curves denote the same frequency choices for $\beta_{M}=1$ and $\beta_{S}=1000$.
Figure 6: Spherical particle, see Fig. 1. Time evolution of negativity $\mathcal{N}(\hat{\rho}_{12})$, computed for $\Omega_{1}=0.54\omega_{p},\Omega_{2}=\omega_{p}$ and the same initial state as in Figs.3 and 4. Solid lines denote the cases $\mu=0.01,0.02,0.04$ and inverse temperatures $\beta_{S}=\beta_{M}=1000$, while dashed lines denote analogous curves computed for $\beta_{M}=1,\beta_{S}=1000$.
Figure 7: Rod-nanodisk configuration, see Fig. 2. Time evolution of negativity $\mathcal{N}(\hat{\rho}_{12})$, computed for $\beta_{M}=\beta_{S}=1000$ and the same initial state as in Figs.3 and 4, for different choices of transition frequencies and dipole momenta.
VI.3 Entanglement Generation
We now address the creation of entanglement starting from separable state preparations. In particular, we confine the initial state of the two emitters to convex mixtures of product states $\sum_{k}p_{k}\,\rho^{k}_{1}\otimes\rho^{k}_{2}$, and focus on entanglement generation in the course of relaxation, keeping in mind that separable states may still display quantum correlations beyond entanglement [58]. To isolate environment-induced effects, we consider the pure product $|\hat{\rho}_{12}(0)=\outerproduct{eg}{eg}$, so that no initial classical or quantum correlations are present. The electromagnetic environment is associated with the dielectric sphere of Fig. 1; where we set the detuning of the two emitter frequencies to zero, i.e., $\Omega_{1}=\Omega_{2}=0.54\omega_{p}$.
In Fig. 8, we first study the evolution of the composite state for fixed coupling strengths and inverse temperatures. The results show that for sufficiently long times, $\mathcal{N}(\hat{\rho}_{12})$ turns positive and starts increasing as a function of time, i.e., a degree of entanglement between the two emitters can be generated as a result of evolution. This property can be traced back to the action of the cross-coupling terms in Eq. (72), which allow for photon-mediated interactions between the two emitters. Similarly to Sect. VI.2, in the case of finite temperature of the medium-assisted reservoir, the system evolves in a separable state for very long times, i.e., in the considered setting, finite temperatures turn out to be detrimental for mediated interactions.
A further dynamical signature of the influence of the cross-coupling interactions along with their spectral features is reported in Fig. 9, where we compare the negativity $\mathcal{N}(\hat{\rho}_{12})$ for time-evolved states of identical emitters with $\Omega_{1}=\Omega_{2}$ while parametrically varying the common transition frequency. When the transition frequencies are tuned near the peak of the cross spectral density $\mathcal{J}_{12}(\omega)$, the value of negativity at moderately long times can be enhanced.
It is also natural to ask whether or not the generation of entanglement states of the emitters at long times decreases with the coupling strength. In principle, increasing the coupling strength of the two emitters can lead to enhancements of these long-time quantum correlations. However, due to the form of Eq.(72), relaxation of the individual emitter states would also increase, as well as decoherence in the state of the composite system. As a consequence, competing effects take place in the course of the dynamics, resulting in a non-monotonic behavior of entanglement with the magnitude of $\mu$. The effects of this competition are evident in Fig. 10, where we plot $\mathcal{N}(\hat{\rho}_{12})$ at low environment temperatures for increasing $\mu$. Notice that the onset time of nonzero negativity increases with $\mu$. This hints at possible steady states that can sustain entanglement between the emitters. Establishing the existence and nature of such steady states would require different numerical approaches [59], which is left for future work.
Figure 8: Spherical particle, see Fig. 1. Time evolution of negativity $\mathcal{N}(\hat{\rho}_{12})$ for the initial composite state $|\hat{\rho}_{12}(0)=\outerproduct{eg}{eg}$, $\Omega_{1}=\Omega_{2}=0.54\omega_{p}$, $\mu=10^{-3}\omega_{p}$, computed for $\beta_{S}=\beta_{M}=1000$ (solid line) and $\beta_{M}=1,\beta_{S}=1000$ (dashed line).
Figure 9: Spherical particle, see Fig. 1. Time evolution of negativity $\mathcal{N}(\hat{\rho}_{12})$ for the initial composite state $|\hat{\rho}_{12}(0)=\outerproduct{eg}{eg}$, $\mu=10^{-3}\omega_{p}$, $\beta_{S}=\beta_{M}=1000$, computed for different values of the emitter frequencies $\Omega_{1}=\Omega_{2}$ in the range $0.30\omega_{p},0.80\omega_{p}$].
Figure 10: Spherical particle, see Fig. 1. Time evolution of negativity $\mathcal{N}(\hat{\rho}_{12})$ for the initial composite state $|\hat{\rho}_{12}(0)=\outerproduct{eg}{eg}$, $\Omega_{1}=\Omega_{2}=0.54\omega_{p}$$\beta_{S}=\beta_{M}=1000$, computed for different dipole strengths $\mu$ in the range $1\cdot 10^{-3}\omega_{p},5\cdot 10^{-3}\omega_{p}]$.
VII Conclusions and outlook
Macroscopic quantum electrodynamics in the context of quantum nanophotonics is a robust framework for describing from an ab initio perspective how a collection of quantum emitters interacts with complex nanophotonic structures, which can be accounted for by using macroscopic Maxwell equations. In this article, we have introduced an extension of the modified Langevin noise approach we recently developed for a single quantum emitter [1], [2] to the case of several quantum emitters. In the modified Langevin noise formalism, the electromagnetic environment consists of two continuous bosonic reservoirs, a medium-assisted bosonic reservoir and a scattering-assisted bosonic reservoir that in general are initially in different quantum states. We have first demonstrated how to define and obtain the medium-assisted and scattering-assisted spectral densities, two distinct matrix-valued functions that fully characterize the interaction of the electromagnetic environment with the quantum emitters. These matrices cannot be evaluated from the knowledge of the dyadic Green function of the dielectric objects at the positions of the quantum emitters. We have shown that they can be effectively computed numerically by surface integral equation formulations of linear classical electromagnetic scattering. We have then introduced the temperature-dependent medium and temperature-dependent scattering-assisted spectral density matrices to describe electromagnetic environments with medium-assisted and scattering-assisted bosonic reservoirs initially in thermal quantum states at different temperatures. Eventually, we have demonstrated that the time evolution of the reduced density operator of the quantum emitters can be obtained by introducing a surrogate bosonic environment, initially in a vacuum quantum state, with an effective spectral density matrix given by the sum of the temperature-dependent medium and temperature-dependent scattering-assisted spectral density matrices. Once the effective spectral density matrix has been calculated, the reduced dynamics of the quantum emitters can be evaluated by using standard approaches for non-Markovian open quantum systems. It is important to note that this framework is applicable not only for deriving precise solutions in quantum nanophotonics, but also as a basis for developing simplified models.
We apply the approach to two quantum emitters that interact with a metallic spherical particle and a metallic nanostructure composed of two rods and a disk. In particular, we found entanglement decay and revivals, as well as entanglement generation. These results underline the utmost importance of the detailed structure of the environment spectral densities in the degradation of quantum correlations. This suggests that the proposed approach can be used to optimize genuinely quantum features of quantum emitters in structured nanophotonics environments.
Future work should focus on extending the proposed approach to the study of molecular polaritons in a dilute ensemble of molecular quantum emitters in optical resonators.
Acknowledgements.
This work was supported by the Italian Ministry of University and Research through the PNRR MUR Project No. PE0000023-NQSTI (C.F. and G.M.) and through the PNRR MUR Project No. CN00000013-ICSC (L.M.C.).
Appendix A Bright electric field operator
The contribution to the electric field operator due to the bright modes can be expressed in terms of the bosonic operators $\hat{C}_{j}$ and $\hat{D}_{j}$. We obtain
$\hat{\mathbf{E}}^{bright}(\mathbf{r})=\sum_{j=1}^{N}\int_{0}^{\infty}d\omega\left[\mathbf{E}_{j}^{(M)}(\mathbf{r},\omega)\hat{C}_{j}(\omega)+h.c.\right]+\\ \sum_{j=1}^{N}\int_{0}^{\infty}d\omega\left[\mathbf{E}_{j}^{(S)}(\mathbf{r},\omega)\hat{D}_{j}(\omega)+h.c.\right]$ (76)
where
$\mathbf{E}_{j}^{(M)}(\mathbf{r},\omega)$ $=\sum_{j=1}^{N}U_{ij}^{*}(\omega)\mathbf{e}_{j}^{(M)}(\mathbf{r},\omega),$ (77a)
$\mathbf{E}_{j}^{(S)}(\mathbf{r},\omega)$ $=\sum_{j=1}^{N}V_{ij}^{*}(\omega)\mathbf{e}_{j}^{(S)}(\mathbf{r},\omega),$ (77b)
and
$\mathbf{e}_{j}^{(M)}(\mathbf{r},\omega)=\int_{V}d^{3}\mathbf{r}^{\prime}\mathcal{G}_{eq}(\mathbf{r},\mathbf{r}^{\prime},\omega)\mathcal{G}^{*T}_{eq}(\mathbf{r}^{\prime},\mathbf{r}_{j},\omega)\mathbb{u}_{j},$ (78a)
$\mathbf{e}_{j}^{(S)}(\mathbf{r},\omega)=\int do_{\mathbb{n}}\sum_{\nu}\mathbf{E}_{\omega\mathbb{n}\nu}(\mathbf{r})\mathbf{E}_{\omega\mathbb{n}\nu}(\mathbf{r}_{j})\mathbb{u}_{j}.$ (78b)
Appendix B Spectral density matrices and power observables in the framework of classical electrodynamics
We here show that the spectral density matrices, introduced in Section IV, are related to power observables of the system of multiple emitters and dielectric objects in the framework of classical electrodynamics. We consider a reciprocal dielectric object in an unbounded space. The object is driven by $N$ electric dipoles, all oscillating harmonically at the angular frequency $\omega$. The $i$-the electric dipole is located at position $\mathbf{r}_{i}$ and has dipole moment $\mathbf{p}_{i}=\mu_{i}\mathbf{u}_{i}$ where $\mu_{i}$, assumed real, is the amplitude and $\mathbf{u}_{i}$ is a unit vector that gives the dipole orientation. The associated electric current density is $\mathbf{J}_{i}=-i\omega\mu_{i}\mathbf{u}_{i}\,\delta\,(\mathbf{r}-\mathbf{r}_{i})$. Because the medium is linear, the total electric field generated by $\mathbf{J}(\mathbf{r})=\sum_{i=1}^{N}\mathbf{J}_{i}(\mathbf{r})$ is the superposition of the fields produced when each dipole acts alone:
$\mathbf{E}(\mathbf{r})=\displaystyle\sum_{j=1}^{N}\mathbf{E}_{j}(\mathbf{r}).$ (79)
The electric field $\mathbf{E}_{j}$, generated by the $j$-th dipole alone, can be expressed via the dyadic Green’s function $\mathcal{G}_{\omega}$ as
$\mathbf{E}_{j}(\mathbf{r})=\frac{\omega^{2}}{\varepsilon_{0}c^{2}}\,\displaystyle\mathcal{G}_{\omega}(\mathbf{r},\mathbf{r}_{j})\cdot(\mu_{j}\mathbf{u}_{j}).$ (80)
Figure 11: Application of the Lorentz reciprocity theorem to two source-field configurations. In Scenario I, only dipole $\mathbf{p}_{i}=\mu_{i}\mathbf{u}_{i}$ at $\mathbf{r}_{i}$ is present, producing fields $(\mathbf{E}_{i},\mathbf{H}_{i})$. In Scenario II, a reference dipole $\mathbf{p}_{\infty}=\mu_{\infty}\mathbf{u}_{\infty}$ is placed at $\mathbf{r}_{\infty}=r_{\infty}\mathbf{n}$ in the far zone of the dielectric object, with $\mathbf{n}\cdot\mathbf{u}_{\infty}=0$, and producing fields $(\mathbf{E}_{\infty},\mathbf{H}_{\infty})$. The origin of the reference system is in the centroid of the dielectric object, that is, the center of the sphere circumscribing the dielectric object with minimum radius $\ell_{c}$.
B.1 Lorentz Reciprocity and a far-field identity
We derive here an identity used later in Sec. B.4. Consider two source–field configurations at the same frequency $\omega$ in the presence of the dielectric object, as depicted in Fig. 11. We choose as the origin of the reference system the centroid of the dielectric object, that is, the center of the sphere that circumscribes the dielectric object with minimum radius $\ell_{c}$. In Scenario I, only a dipole $\mathbf{p}_{i}=\mu_{i}\mathbf{u}_{i}$ is present at $\mathbf{r}_{i}$ and produces the electromagnetic field $\mathbf{E}_{i}(\mathbf{r})$. In Scenario II, a reference dipole $\mathbf{p}_{\infty}=\mu_{\infty}\mathbf{u}_{\infty}$ produces the electric field $\mathbf{E}_{\infty}(\mathbf{r})$. It is placed at $\mathbf{r}_{\infty}=r_{\infty}\mathbf{n}$ in the Fraunhofer zone (far zone) of the dielectric object, i.e. $r_{\infty}\gg k_{\omega}\ell_{c}^{2}/\pi$, where $\mathbf{u}_{\infty}$ and $\mathbf{n}$ are unit vectors with $\mathbf{n}\cdot\mathbf{u}_{\infty}=0$; the associated current density is $\mathbf{J}_{\infty}=-i\omega\,\mu_{\infty}\,\mathbf{u}_{\infty}\,\delta(\mathbf{r}-\mathbf{r}_{\infty})$. The Lorentz reciprocity theorem yields [60]
$\int_{\mathbb{R}^{3}}\mathbf{E}_{i}(\mathbf{r})\cdot\mathbf{J}_{\infty}(\mathbf{r})\,dV=\int_{\mathbb{R}^{3}}\mathbf{E}_{\infty}(\mathbf{r})\cdot\mathbf{J}_{i}(\mathbf{r})\,dV,$ (81)
which implies
$\mu_{\infty}\,\mathbf{E}_{i}(\mathbf{r}_{\infty})\cdot\mathbf{u}_{\infty}=\mu_{i}\,\mathbf{E}_{\infty}(\mathbf{r}_{i})\cdot\mathbf{u}_{i}.$ (82)
At position $\mathbf{r}_{\infty}$ in the Fraunhofer zone of the dielectric objects, we have in Scenario I
$\mathbf{E}_{i}(\mathbf{r}_{\infty}){\approx}\frac{\mathbf{A}_{i}(\mathbf{n})}{r_{\infty}}\,e^{ik_{\omega}r_{\infty}}$ (83)
where $\mathbf{A}_{i}(\mathbf{n})$ is the far-field pattern of $\mathbf{E}_{i}$ with $\mathbf{n}\cdot\mathbf{A}_{i}(\mathbf{n})=0$. In Scenario II, the incident electric field generated at the point $\mathbf{r}_{i}$, generated by the remote dipole alone, is
$\mathbf{E}_{\infty}^{inc}(\mathbf{r}_{i})\approx\frac{k_{\omega}^{2}}{4\pi\varepsilon_{0}}\,\frac{\mu_{\infty}}{r_{\infty}}\exp\left(ik_{\omega}\,\mathbf{n}\cdot\mathbf{r}_{i}\right)\exp\left(-ik_{\omega}r_{\infty}\right)\mathbf{u}_{\infty}$ (84)
as $\mathbf{r}_{i}$ lies in the Fraunhofer zone of the electromagnetic field generated by the dipole $\mathbb{p}_{\infty}$. Thus, in Scenario II, the dielectric object experiences an incident field that is a transverse plane wave propagating along $\mathbf{n}$, polarized along $\mathbf{u}_{\infty}$ and with unit amplitude provided that $\mu_{\infty}=4\pi\varepsilon_{0}\,r_{\infty}/{k_{\omega}^{2}}$. According to Eqs. 7-8, we denote by $\mathbf{F}_{\omega\mathbf{n}\mathbf{u}_{\infty}}(\mathbf{r}_{i})$ the total electric field at $\mathbf{r}_{i}$ scattered by the dielectric objects when excited by a plane wave of unit intensity propagating along $\mathbf{n}$ and polarized along $\mathbf{u}_{\infty}$, hence
$\mathbf{E}_{\infty}(\mathbf{r}_{i})=\mathbf{F}_{\omega\mathbf{n}\mathbf{u}_{\infty}}(\mathbf{r}_{i}).$ (85)
This equation is valid up to the overall phase factor $\exp\left({-ik_{\omega}r_{\infty}}\right)$ that cancels in the power observables. Using this result in (82) together with (83) we obtain the projection identity
$\mathbf{A}_{i}(\mathbf{n})\cdot\mathbf{u}_{\infty}=\frac{k_{\omega}^{2}}{4\pi\varepsilon_{0}}\,\mu_{i}\,\mathbf{F}_{\omega\mathbf{n}\mathbf{u}_{\infty}}(\mathbf{r}_{i})\cdot\mathbf{u}_{i},$ (86)
In the scattering-mode normalization introduced in Eq. (9), Eq. 86 becomes:
$\mathbf{A}_{i}(\mathbf{n})\cdot\mathbf{u}_{\infty}=\sqrt{\frac{\pi\omega}{\hbar\varepsilon_{0}c}}\,\mu_{i}\,[\mathbf{E}_{\omega\mathbf{n}\mathbf{u}_{\infty}}(\mathbf{r}_{i})\cdot\mathbf{u}_{i}].$ (87)
This relation states that the $\mathbf{u}_{\infty}$-polarized far-field amplitude in direction $\mathbf{n}$ generated by the dipole at $\mathbb{p}_{i}$ is proportional to the projection of the normalized scattering mode $\mathbf{E}_{\omega\mathbf{n}\mathbf{u}_{\infty}}$ at the dipole location onto the dipole orientation $\mathbf{u}_{i}$. In other words, it links the coupling between the dipole and the scattering mode at $\mathbf{r}_{i}$ to the corresponding far-field pattern produced when the object is driven by a dipole $\mu_{i}\mathbf{u}_{i}$ at $\mathbf{r}_{i}$.
B.2 Time-averaged power emitted by the dipoles
The time-averaged power emitted by the $N$ dipoles is
$\mathscr{P}^{em}=-\frac{1}{2}\text{Re}\int_{\mathbb{R}^{3}}d^{3}\mathbf{r}\,\mathbf{J}^{*}\cdot\mathbf{E}=\sum_{i,j=1}^{N}\mathscr{P}_{ij}^{em},$ (88)
where
$\mathscr{P}^{em}_{ij}=-\frac{1}{2}\text{Re}\int_{\mathbb{R}^{3}}d^{3}\mathbf{r}\,\mathbf{J}_{i}^{*}\cdot\mathbf{E}_{j}.$ (89)
Using 80 in 89, one finds the following
$\mathscr{P}_{ij}^{em}=\frac{\mu_{0}}{2}\omega^{3}(\mu_{i}\mathbf{u}_{i})\cdot\text{Im}\left[\mathcal{G}_{\omega}(\mathbf{r}_{i},\mathbf{r}_{j})\right]\cdot(\mu_{j}\mathbf{u}_{j}).$ (90)
A comparison of the above equation with Eq. (54) yields
$\mathscr{P}_{ij}^{em}=\frac{\hbar\omega}{4}\Gamma_{ij}.$ (91)
B.3 Time-averaged power absorbed by the dielectric object
The time-averaged power absorbed by the dielectric objects is
$\mathscr{P}^{abs}=\frac{\omega}{2}\varepsilon_{0}\int_{V}d^{3}\mathbf{r}\,\text{Im}[\varepsilon_{\omega}(\mathbf{r})]\mathbf{E}(\mathbf{r})\cdot\mathbf{E}^{*}(\mathbf{r})\\ =\sum_{i,j=1}^{N}\mathscr{P}^{abs}_{ij}$ (92)
with
$\mathscr{P}^{abs}_{ij}=\frac{\omega}{2}\varepsilon_{0}\int_{V}d^{3}\mathbf{r}\,\text{Im}[\varepsilon_{\omega}(\mathbf{r})]\mathbf{E}_{i}(\mathbf{r})\cdot\mathbf{E}^{*}_{j}(\mathbf{r}).$ (93)
Using 80 and the reciprocity for the dyadic Green function we get:
$\mathscr{P}^{abs}_{ij}=\frac{1}{2}\varepsilon_{0}\mu_{0}^{2}\omega^{5}\mu_{i}\mu_{j}\times\\ \int_{V}d^{3}\text{Im}[\varepsilon_{\omega}(\mathbf{r})]\mathbf{u}_{i}\cdot\mathcal{G}_{\omega}(\mathbf{r}_{i},\mathbf{r})\mathcal{G}_{\omega}^{*T}(\mathbf{r}_{j},\mathbf{r})\cdot\mathbf{u}_{j}$ (94)
where we have used the property $\mathcal{G}_{\omega}(\mathbf{r},\mathbf{r}_{j})=\mathcal{G}_{\omega}^{T}(\mathbf{r}_{j},\mathbf{r})$. Using the relation between $\mathcal{G}_{e}$ and $\mathcal{G}_{\omega}$ in Eq. 5, we get:
$\mathscr{P}^{abs}_{ij}=\frac{\pi}{2}{\hbar\omega}\frac{\mu_{i}\mu_{j}}{\hbar^{2}}\times\\ \int_{V}d^{3}\mathbf{r}\,\mathbf{u}_{i}\cdot\mathcal{G}_{e}(\mathbf{r}_{i},\mathbf{r},\omega)\mathcal{G}_{e}^{*T}(\mathbf{r}_{j},\mathbf{r},\omega)\cdot\mathbf{u}_{j}=\frac{\pi}{2}{\hbar\omega}\mathcal{J}^{(M)}_{ij}$ (95)
where $\mathcal{J}^{(M)}_{ij}$ is defined in Eq. (52a).
B.4 Averaged Power Radiated to infinity
The time-averaged power radiated through a sphere $S_{\infty}$ at infinity is
$\mathscr{P}^{rad}=\frac{1}{4}\int_{S_{\infty}}\,do_{\mathbf{n}}r_{\infty}^{2}\left(\mathbf{E}\times\mathbf{H}^{*}+\mathbf{E}^{*}\times\mathbf{H}\right)\cdot\mathbf{n}=\\ \sum_{i,j=1}^{N}\mathscr{P}^{rad}_{ij}$ (96)
with
$\mathscr{P}^{rad}_{ij}=\oint_{S_{\infty}}\,do_{\mathbf{n}}r_{\infty}^{2}\mathbf{S}_{ij}(r_{\infty},\mathbf{n})\cdot\mathbf{n},$ (97)
and the symmetrized mutual Poynting vector is defined as:
$\mathbf{S}_{ij}=\frac{1}{4}\left(\mathbf{E}_{i}\times\mathbf{H}_{j}^{*}+\mathbf{E}_{j}^{*}\times\mathbf{H}_{i}\right).$ (98)
In the far zone, we have
[…]
The Poynting vector 98 becomes
$\mathbf{S}_{ij}=\frac{1}{2\zeta_{0}}\frac{1}{r_{\infty}^{2}}\left[\mathbf{A}_{i}(\mathbf{n})\cdot\mathbf{A}_{j}^{*}(\mathbf{n})\right]\,\mathbf{n}.$ (100)
By replacing this expression in 97 we get:
$\mathscr{P}^{rad}_{ij}=\frac{1}{2\zeta_{0}}\oint_{S_{\infty}}\,do_{\mathbf{n}}\,\displaystyle\sum_{\nu}(\mathbf{A}_{i}\cdot\bm{\nu})(\mathbf{A}_{j}^{*}\cdot\bm{\nu})$ (101)
Using relation 87 one finds
$\mathscr{P}^{rad}_{ij}=\frac{\pi}{2}{\hbar\omega}\frac{\mu_{i}\mu_{j}}{\hbar^{2}}\times\\ \oint_{S_{\infty}}\,do_{\mathbf{n}}\,\sum_{\nu}\mathbf{u}_{i}\cdot\mathbf{E}_{\omega\mathbf{n}\nu}(\mathbf{r}_{i})\mathbf{E}^{*}_{\omega\mathbf{n}\nu}(\mathbf{r}_{j})\cdot\mathbf{u}_{j}=\frac{\pi}{2}{\hbar\omega}\mathrm{\mathcal{J}}_{ij}^{(S)}(\omega)$ (102)
where $\mathcal{J}^{(S)}_{ij}$ is defined in Eq. (52b).
In summary, the spectral-density matrices $\mathcal{J}^{(S)}_{ij}$ and $\mathcal{J}^{(M)}_{ij}$ can be computed purely within classical electrodynamics by evaluating the pairwise radiated and absorbed powers $\mathscr{P}^{rad}_{ij}$ and $\mathscr{P}^{abs}_{ij}$.
B.5 Mutual Poynting theorem
Direct evaluation of Eq. (95) requires a volume integral over $V$, and hence a volumetric mesh. This can be avoided by recasting the calculation as a surface integral.
Let $\Omega$ be any volume whose boundary $\partial\Omega$ encloses the dielectric body but excludes the sources $i$ and $j$ (Fig. 12). Using Maxwell’s equations in $\Omega$ and taking the divergence of $\mathbf{S}_{ij}$ in (98) yields the mutual Poynting theorem
Figure 12: Proof of the mutual Poynting theorem. $\Omega$ is any volume whose boundary $\partial\Omega$ encloses the dielectric body but excludes the sources $i$ and $j$.
$\mathscr{P}_{ij}^{abs}=\frac{\omega}{2}\varepsilon_{0}\int_{V}d^{3}\mathbf{r}\,\text{Im}[\varepsilon_{\omega}(\mathbf{r})]\mathbf{E}_{i}(\mathbf{r})\cdot\mathbf{E}_{j}^{*}(\mathbf{r})\\ =-\oint_{\partial\Omega}d^{2}\mathbf{r}\,\mathbf{S}_{ij}\,\cdot\mathbf{m}$ (103)
where $\mathbf{m}$ denotes the outward-pointing unit normal vector on $\partial\Omega$. Equation (103) allows one to rewrite the volume integral entering $\mathcal{J}^{(M)}_{ij}$ as a surface integral over $\Sigma$, which is advantageous for numerical evaluation with a surface-integral-equation solver.
Appendix C Correlators
In this Section we compute the elements $C_{ij}^{(M)}(t)$ of the correlator matrix of the medium-assisted reservoir; a similar result holds for the elements of the correlator matrix of the scattering-assisted reservoir.
We consider the following scenario: (a) the initial quantum state of the electromagnetic environment is a product state, that is, $\hat{\rho}_{E}(0)=\hat{\rho}_{E}^{(M)}(0)\otimes\hat{\rho}_{E}^{(S)}$, where $\hat{\rho}_{E}^{(M)}(0)$ and $\hat{\rho}_{E}^{(S)}(0)$ are the initial density operators of the medium-assisted reservoir and the scattering-assisted reservoir, respectively; (b) the two reservoirs are initially in thermal states with temperatures $T_{0}^{(M)}$ and $T_{0}^{(S)}$.
The correlation function $C^{(M)}_{ij}(t)$ is given by
$C^{(M)}_{ij}(t)=\langle\hat{F}^{(M)}_{i}(t)\hat{F}^{(M)}_{j}(0)\rangle$ (104)
where $\hat{{F}}^{(M)}_{i}(t)$ is the interaction operator of the medium-assisted reservoir in the interaction picture,
$\hat{{F}}^{(M)}_{i}(t)=\hat{U}^{\dagger}_{E}(t)\hat{F}^{(M)}_{i}\hat{U}_{E}(t),$ (105)
$\hat{U}_{E}(t)=\exp(-i\hat{H}_{E}^{bright}t/\hbar)$ is the free evolution operator of the bright bosonic modes of the electromagnetic environment, and $\langle\cdot\rangle=Tr_{E_{M}}[\cdot\,\rho_{E}^{(M)}(0)]$. Starting with 105, and using Eqs. 32, 46, and the commutation relations 30 we obtain
$\hat{F}_{i}^{(M)}(t)=\sum_{j=1}^{N}\int_{0}^{\infty}d\omega[G_{ij}^{(M)}(\omega)\hat{C}_{j}(\omega)e^{-i\omega t}+h.c.].$ (106)
Substituting 106 into 104, and using again 30 we find
$C_{ij}^{(M)}(t)=\int_{0}^{\infty}d\omega\sum_{h=1}^{N}\left[G^{(M)}_{in}(\omega)G^{(M)*}_{jn}(\omega)\langle\hat{C}_{n}(\omega)\hat{C}^{\dagger}_{n}(\omega)\rangle e^{-i\omega t}+G^{(M)*}_{in}(\omega)G^{(M)}_{jn}(\omega)\langle\hat{C}^{\dagger}_{n}(\omega)\hat{C}_{n}(\omega)\rangle e^{+i\omega t}\right],$ (107)
because for a thermal state $\langle\hat{C}_{m}(\omega)\hat{C}_{n}(\omega)\rangle=\langle\hat{C}_{m}^{\dagger}(\omega)\hat{C}_{n}^{\dagger}(\omega)\rangle=0,$ $\langle\hat{C}_{m}(\omega)\hat{C}_{n}^{\dagger}(\omega)\rangle=\delta_{mn}\big(1+n_{\omega}^{(M)}\big)$, $\langle\hat{C}_{m}^{\dagger}(\omega)\hat{C}_{n}(\omega)\rangle=\delta_{mn}\,n_{\omega}^{(M)},$ where
$n_{\omega}^{(M)}=\langle\hat{C}_{n}^{\dagger}(\omega)\hat{C}_{n}(\omega)\rangle=\frac{1}{e^{\beta_{M}\hbar\omega}-1},$ (108)
and $\beta_{M}=1/(k_{B}T_{0}^{(M)})$. Exploiting that
$\langle\hat{C}_{n}(\omega)\hat{C}_{n}^{\dagger}(\omega)\rangle=1+\langle\hat{C}_{n}^{\dagger}(\omega)\hat{C}_{n}(\omega)\rangle,$ (109)
$G_{jn}^{(M)*}(\omega)=G_{nj}^{(M)}(\omega)$ and $\sum^{N}_{n=1}G^{(M)}_{in}(\omega)G^{(M)}_{nj}(\omega)=M_{ij}$, Eq. 107 becomes
$C_{ij}^{(M)}(t)=\int_{0}^{\infty}\mathrm{d}\omega[(1+n_{\omega}^{(M)}){M}_{ij}(\omega)e^{-i\omega t}+n_{\omega}^{(M)}{M}_{ij}^{*}(\omega)e^{i\omega t}].$ (110)
Appendix D Surface Integral Equation Method
Surface–integral–equation (SIE) formulations [61] are well suited for the classical evaluation of spectral-density matrices when the dielectric object is piecewise homogeneous: unknowns live only on object boundaries and the radiation condition at infinity is inherently satisfied.
We consider a single dielectric body that occupies a bounded volume $V$ with a boundary $\partial V$. The interior (“++”) medium is homogeneous, with $\varepsilon^{+}(\omega)=\varepsilon_{0}\,\varepsilon_{\omega}$ and $\mu^{+}=\mu_{0}$. The exterior (“-”) medium is vacuum, with $\varepsilon^{-}=\varepsilon_{0}$ and $\mu^{-}=\mu_{0}$. Let $k_{\omega}^{\pm}=\omega\sqrt{\mu^{\pm}\varepsilon^{\pm}}$ and $\zeta^{\pm}=\sqrt{\mu^{\pm}/\varepsilon^{\pm}}$ denote, respectively, the wavenumbers and wave impedances. The object is illuminated by a time–harmonic field, $\mathrm{Re}\{\mathbf{E}_{\mathrm{inc}}(\mathbf{r})\,e^{-i\omega t}\}$, $\mathrm{Re}\{\mathbf{H}_{\mathrm{inc}}(\mathbf{r})\,e^{-i\omega t}\}$.
Following Poggio–Miller–Chang–Harrington–Wu–Tsai (PMCHWT) [62, 63, 64], the equivalent electric and magnetic surface currents $\mathbf{j}_{e}(\mathbf{r})$ and $\mathbf{j}_{m}(\mathbf{r})$, defined on $\partial V$, solve:
$\bm{\mathcal{Z}}\,\mathbf{j}=\mathbf{v},$ (111)
with block operators
[…]
$\mathbf{j}=\left[\mathbf{j}_{e},\mathbf{j}_{m}\right]^{T}$, $\mathbf{v}=\left[\mathbf{e}_{0},\mathbf{h}_{0}\right]^{T}$, $\mathbf{e}_{0}=\left.-\mathbf{n}\times\mathbf{n}\times{\bf{E}}_{inc}\right|_{\partial\Omega}$, and $\mathbf{h}_{0}=\left.-\mathbf{n}\times\mathbf{n}\times{\bf{H}}_{inc}\right|_{\partial\Omega}.$ The EFIE/MFIE boundary operators $\mathcal{T}_{\pm}$ and $\mathcal{K}_{\pm}$ act on a tangential test function $\mathbf{w}$ as:
$\bm{\mathcal{K}}_{\pm}$ $\left\{{\bf w}\right\}\left({\bf r}\right)=\mathbf{n}\times\mathbf{n}\times\int_{\partial\Omega}{\bf w}\left({\bf r}^{\prime}\right)\times\nabla^{\prime}g^{\pm}\left({\bf r}-{\bf r}^{\prime}\right)dS^{\prime},$ (113a)
$\bm{\mathcal{T}}_{\pm}$ $\left\{{\bf w}\right\}\left({\bf r}\right)=-ik^{\pm}_{\omega}\mathbf{n}\times\mathbf{n}\times\int_{\partial\Omega}g^{\pm}\left({\bf r}-{\bf r}^{\prime}\right){\bf w}\left({\bf r}^{\prime}\right)dS^{\prime}$
$-\frac{1}{ik^{\pm}_{\omega}}\mathbf{n}\times\mathbf{n}\times\int_{\partial\Omega}\nabla^{\prime}g^{\pm}\left({\bf r}-{\bf r}^{\prime}\right)\nabla^{\prime}_{S}\cdot{\bf w}\left({\bf r}^{\prime}\right)dS^{\prime},$ (113b)
where
$g^{\pm}\left({\bf r}-{\bf r}^{\prime}\right)=\frac{e^{ik^{\pm}_{\omega}\left|{\bf r}-{\bf r}^{\prime}\right|}}{4\pi\left|{\bf r}-{\bf r}^{\prime}\right|}.$ (114)
is the homogeneous-space Green function.
D.1 Finite-dimensional representation
Figure 13: Illustration of the RWG basis function associated to the $p$ th edge and defined on the triangle pair $T_{p}^{+},T_{p}^{-}$ with centroids $\mathbf{g}_{p}^{\pm}$. In $T_{p}^{+}$and $T_{p}^{-}$ the vector field $\mathbf{f}_{p}$ is proportional to the vector ( $\mathbf{r}-\mathbf{v}_{p}^{+}$) and ( $\mathbf{v}_{p}^{-}-\mathbf{r}$ ), respectively. To each RWG basis function is associated a dipole moment $\mathbf{p}_{p}$ centered in the midpoint $\mathbf{c}_{p}=(\mathbf{g}_{p}^{+}+\mathbf{g}_{p}^{-})/2$.
We discretize $\partial V$ with a conforming triangular mesh $\mathcal{M}$ having $N_{e}$ interior edges. For edge $p$, let $\ell_{p}$ be its length; $T_{p}^{\pm}$ the adjacent triangles with areas $A_{p}^{\pm}$; $\mathbf{v}_{p}^{\pm}$ the vertices opposite to the common edge; and $\mathbf{g}_{p}^{\pm}$ their centroids. Define the centroid offset $\mathbf{d}_{p}=\mathbf{g}_{p}^{+}-\mathbf{g}_{p}^{-}$ and the midpoint $\mathbf{c}_{p}=(\mathbf{g}_{p}^{+}+\mathbf{g}_{p}^{-})/2$. The Rao–Wilton–Glisson (RWG) basis function $\mathbf{f}_{p}$ is the piecewise linear tangential field
$\mathbf{f}_{p}(\mathbf{r})=\frac{\ell_{p}}{2}\times\begin{cases}\left(\mathbf{r}-\mathbf{v}_{p}^{+}\right)/A_{p}^{+}&\mathbf{r}\in T_{p}^{+}\\ \left(\mathbf{v}_{p}^{-}-\mathbf{r}\right)/A_{p}^{-}&\mathbf{r}\in T_{p}^{-}\\ 0&\text{elsewhere}\end{cases}$
whose support is $S_{p}=T_{p}^{+}\cup T_{p}^{-}$. Its surface divergence is constant over each triangle. Thus, because of the continuity equation, the RWG basis function $\mathbf{f}_{p}$ is associated on the triangle $T_{p}^{\pm}$ to a total electric charge $Q_{p}^{\pm}=\pm\ell_{p}/(i\omega)$, which can be thought as localized in its centroid $\mathbf{g}_{p}^{+}$, and a total magnetic charge $Q_{m}^{\pm}=\pm\ell_{p}/(i\omega\mu_{0})$ localized in its centroid $\mathbf{g}_{p}^{-}$.
We expand the unknown currents in the RWG basis functions $\{\mathbf{f}_{p}\}_{p=1}^{N_{e}}$:
$\mathbf{j}_{e}(\mathbf{r})\approx\sum_{p=1}^{N_{e}}\alpha_{p}\mathbf{f}_{p}(\mathbf{r}),\qquad\mathbf{j}_{m}(\mathbf{r})\approx\sum_{p=1}^{N_{e}}\beta_{p}\mathbf{f}_{p}(\mathbf{r}).$ (115)
Galerkin testing with the same basis yields the finite–dimensional counterpart of (111):
$\mathbf{Z}\,\mathbf{J}=\mathbf{V}$ (116)
with
$\mathbf{Z}=\begin{bmatrix}\zeta^{-}T_{-}+\zeta^{+}T_{+}&K_{-}+K_{+}\\ -(K_{-}+K_{+})&\frac{1}{\zeta^{-}}T_{-}+\frac{1}{\zeta^{+}}T_{+}\end{bmatrix}$, (117)
$T_{\pm}]_{ij}=\langle\mathbf{f}_{i}|\mathcal{T}_{\pm}|\mathbf{f}_{j}\rangle,\hskip 21.52771pt[K_{\pm}]_{ij}=\langle\mathbf{f}_{i}|\mathcal{K}_{\pm}|\mathbf{f}_{j}\rangle,$ (118)
$\mathbf{J}=[\mathbf{J}_{e},\mathbf{J}_{m}]^{T}$, $\mathbf{V}=\left[\mathbf{E}_{0},\mathbf{H}_{0}\right]^{T}$, $\mathbf{J}_{e}=[\alpha_{1},\ldots,\alpha_{N_{e}}]^{\intercal}$, $\mathbf{J}_{m}=[\beta_{1},\ldots,\beta_{N_{e}}]^{\intercal}$, $\mathbf{E}_{0}]_{i}=\langle\mathbf{f}_{i}|\mathbf{e}_{0}\rangle$, $\mathbf{H}_{0}]_{i}=\langle\mathbf{f}_{i}|\mathbf{h}_{0}\rangle$, and
$\langle\mathbf{u}|\mathbf{v}\rangle=\int_{\partial\Omega}d^{2}\mathbf{r}\mathbf{u}^{*}(\mathbf{r})\cdot\mathbf{v}(\mathbf{r}).$ (119)
To evaluate pairwise absorbed and radiated powers $\mathscr{P}^{\mathrm{abs}}_{ij}$ and $\mathscr{P}^{\mathrm{rad}}_{ij}$, many right–hand sides are required (one per impressed dipole). When $N$ is large, it is efficient to compute a single LU factorization of $\mathbf{Z}$ and reuse it for all right–hand sides by matrix-vector multiplications. We denote $\{\alpha_{p}^{(i)}\}_{p\in N_{e}}$ $\{\beta_{p}^{(i)}\}_{p\in N_{e}}$ the expansion coefficients for the $i$-th impressed dipole excitation.
D.2 Time-averaged power absorbed by the dielectric object
We evaluate $\mathscr{P}^{\mathrm{abs}}_{ij}$ via the mutual Poynting theorem, Eq. (103). A natural choice for $\partial\Omega$ is the particle boundary $\partial V$, where the fields can be obtained directly from the solved equivalent surface currents (and the associated surface charges). However, for irregularly shaped bodies the surface fields on $\partial V$ may be large and rapidly varying, which can slow the convergence of the surface integral. In such cases it is advantageous to choose a smooth enclosing surface $\partial\Omega$ , e.g. a spherical/spheroidal surface, that surrounds $V$ but lies a distance away from it, provided $\partial\Omega$ encloses $V$ and excludes the impressed sources.
D.3 Average Powers radiated to infinity
To compute the $\mathscr{P}_{ij}^{rad}$ we solve the SIE scattering problem for two excitations: an impressed dipole $\mathbf{p}_{i}$ at $\mathbf{r}_{i}$ and, separately, $\mathbf{p}_{j}$ at $\mathbf{r}_{j}$. Denote the corresponding RWG expansion coefficients by $\{\alpha_{s}^{(i)},\beta_{s}^{(i)}\}$ and $\{\alpha_{t}^{(j)},\beta_{t}^{(j)}\}$. Then, a brute-force but expensive procedure would evaluate the far-field surface integral in Eq. (101) by forming the radiation from the equivalent surface currents at all quadrature points on $S_{\infty}$ for both $\mathbf{p}_{j}$ and $\mathbf{p}_{j}$ excitations, which is an expensive procedure.
Instead, $\mathscr{P}_{ij}^{\mathrm{rad}}$ can be obtained analytically from the RWG expansion coefficients. As noted in Sec. D.1, the RWG basis function $\mathbf{f}_{p}$ induces equal and opposite electric or magnetic charges on $T_{p}^{\pm}$, thus a single coefficient $\alpha_{p}$ or $\beta_{p}$ corresponds to equivalent electric or magnetic dipole moments (centered at $\mathbf{c}_{p}$, oriented along $\mathbf{d}_{p}$)
$\mathbf{p}_{p}^{(i)}=\alpha_{p}^{(i)}\frac{\ell_{p}}{i\omega}\mathbf{d}_{p},\quad\mathbf{m}_{p}^{(i)}=\beta_{p}^{(i)}\frac{\ell_{p}}{i\omega\mu_{0}}\mathbf{d}_{p},$ (120)
with $i=1,\ldots,N_{e}$. For easy of reference, set $\mathbf{p}_{0}^{(i)}=\mathbf{p}_{i}$ and $\mathbf{c}_{0}=\mathbf{r}_{i}$ (There is no impressed magnetic dipole).
The mutual radiated power is then the sum of free–space pairwise contributions between the effective dipoles of the two solutions
$\mathscr{P}_{ij}^{rad}=\displaystyle\sum_{s,t=0}^{N_{e}}\mathcal{P}_{ee}(\mathbf{p}_{s}^{(i)},\mathbf{p}_{t}^{(j)})+\displaystyle\sum_{s,t=1}^{N_{e}}\mathcal{P}_{mm}(\mathbf{m}_{s}^{(i)},\mathbf{m}_{t}^{(j)})\\ +\displaystyle\sum_{s,t=1}^{N_{e}}\left[\mathcal{P}_{em}(\mathbf{p}_{s}^{(i)},\mathbf{m}_{t}^{(j)})+\mathcal{P}_{em}(\mathbf{m}_{s}^{(i)},\mathbf{p}_{t}^{(j)})\right].$ (121)
The first sum takes into account the pairwise interactions of the electric dipoles, including the impressed ones that occur for $s=0$ and $t=0$:
$\mathcal{P}_{ee}(\mathbf{p}_{s},\mathbf{p}_{t})=\frac{k_{\omega}^{4}}{8\pi\varepsilon_{0}^{2}\zeta_{0}}\times\\ \left\{\mathbf{p}_{s}\cdot\mathbf{p}_{t}^{*}f(k_{\omega}r_{st})-\left(\mathbf{p}_{s}\cdot\hat{\mathbf{r}}_{st}\right)\left(\mathbf{p}_{t}^{*}\cdot\hat{\mathbf{r}}_{st}\right)g(k_{\omega}r_{st})\right\}.$ (122)
The second sum takes into account the pairwise interactions of magnetic dipoles, where $\mathcal{P}_{mm}$ is directly related to $\mathcal{P}_{ee}$ by the duality property of the electromagnetic field:
$\mathcal{P}_{mm}^{ij}(\mathbf{m}_{s},\mathbf{m}_{t})=\frac{1}{c^{2}}\mathcal{P}_{ee}^{ij}(\mathbf{m}_{s},\mathbf{m}_{t}).$ (123)
The third sum takes into account the pairwise interactions of the electric and magnetic dipoles and viceversa with
$\mathcal{P}_{em}^{ij}(\mathbf{p}_{s},\mathbf{m}_{t})=i\frac{1}{c}\frac{k_{\omega}^{4}}{8\pi\varepsilon_{0}^{2}\zeta_{0}}\\ \mathbf{p}_{s}\times\mathbf{m}^{*}_{t}\cdot\hat{\mathbf{r}}_{st}\,h(k_{\omega}r_{st})$ (124)
where ${\mathbf{r}}_{st}=\mathbf{c}_{s}-\mathbf{c}_{t}$, ${r}_{st}=\left|\mathbf{r}_{st}\right|$, $\hat{\mathbf{r}}_{st}={\mathbf{r}}_{st}/r_{st}$, and
$\left\{\begin{aligned} f(x)&=\frac{\sin x}{x}-\,\frac{\sin x-x\cos x}{x^{3}},\\ g(x)&=\frac{\sin x}{x}-3\,\frac{\sin x-x\cos x}{x^{3}},\\ h(x)&=\frac{\sin(x)-x\cos(x)}{x^{2}},\end{aligned}\right$. (125)
$f(x\rightarrow 0)={2}/{3}$, $g(x\rightarrow 0)=0$, and $h(x\rightarrow 0)=0$.
Appendix E Lorentz Reciprocity and a far-field identity (alternative proof)
We derive here an identity used in Sec. B.4. Consider two source–field configurations at the same frequency $\omega$ in the presence of the dielectric object, as depicted in Fig. 11. In Scenario I, only a dipole $\mathbf{p}_{i}=\mu_{i}\mathbf{u}_{i}$ is present at $\mathbf{r}_{i}$ and produces the electromagnetic field $(\mathbf{E}_{i},\mathbf{H}_{i})$. At position $\mathbf{r}_{\infty}$ at infinity
$\mathbf{E}_{i}(\mathbf{r}_{\infty})$ $=\frac{e^{ik_{\omega}r_{\infty}}}{r_{\infty}}\mathbf{A}_{i}(\mathbf{n})+\mathcal{O}(r_{\infty}^{-2})\,$ (126a)
$\mathbf{H}_{i}(\mathbf{r}_{\infty})$ $=\frac{e^{ik_{\omega}r_{\infty}}}{r_{\infty}}\frac{\mathbf{n}\times\mathbf{A}_{i}(\mathbf{n})}{\zeta_{0}}+\mathcal{O}(r_{\infty}^{-2})$ (126b)
where $\mathbf{A}_{i}(\mathbf{n})$ is the far-field pattern of $\mathbf{E}_{i}$ with $\mathbf{n}\cdot\mathbf{A}_{i}(\mathbf{n})=0$.
In Scenario II, the object is illuminated by a unit-amplitude plane wave with propagation direction $\mathbf{n}_{\infty}$ and polarization $\mathbf{u}_{\infty}$:
$\mathbf{E}_{2}^{inc}$ $=\exp\left(ik_{\omega}\mathbf{n}_{\infty}\cdot\mathbf{r}\right)\mathbf{u}_{\infty}$ (127a)
$\mathbf{H}_{2}^{inc}$ $=\frac{1}{\zeta_{0}}\mathbf{n}_{\infty}\times\mathbf{u}_{\infty}\exp\left(ik_{\omega}\mathbf{n}_{\infty}\cdot\mathbf{r}\right)$ (127b)
By linearity, the total fields are
$\mathbf{E}_{2}=\mathbf{E}_{2}^{inc}+\mathbf{E}_{2}^{sca};\qquad\mathbf{H}_{2}=\mathbf{H}_{2}^{inc}+\mathbf{H}_{2}^{sca}$ (128)
with scattered fields $\mathbf{E}_{2}^{sca}$ and $\mathbf{H}_{2}^{sca}$ satisfying the Silver–Müller condition,
$\mathbf{E}^{sca}_{2}(\mathbf{n},r_{\infty})$ $=\frac{e^{ik_{\omega}r_{\infty}}}{r_{\infty}}\,\mathbf{E}_{\infty}({\mathbf{n}})+\mathcal{O}(r_{\infty}^{-2}),$ (129a)
$\mathbf{H}_{2}^{sca}(\mathbf{n},r_{\infty})$ $=\frac{e^{ik_{\omega}r_{\infty}}}{r_{\infty}}\,\frac{\mathbf{n}\times\mathbf{E}_{\infty}({\mathbf{n}})}{\zeta_{0}}+\mathcal{O}(r_{\infty}^{-2}).$ (129b)
Note that the total electric field $\mathbf{E}_{2}$ coincides with $\mathbf{F}_{\omega\mathbf{n}_{\infty}\mathbf{u}_{\infty}}$ as defined in Eqs. (7)–(8).
The Lorentz reciprocity theorem gives
$\oint_{S_{\infty}}d^{2}\mathbf{r}\,\left(\mathbf{E}_{i}\times\mathbf{H}_{2}-\mathbf{E}_{2}\times\mathbf{H}_{i}\right)\cdot\mathbf{n}=\int_{\mathbb{R}^{3}}d^{3}\;\mathbf{r}\mathbf{E}_{2}\cdot\mathbf{J}_{i}$ (130)
where $S_{\infty}$ is a spherical surface at infinity and $\mathbf{J}_{i}$ is the current density of the dipole in Scenario I $\mathbf{J}_{i}=-i\omega\mu_{i}\mathbf{u}_{i}\,\delta(\mathbf{r}-\mathbf{r}_{i})$. Using the far-field expansions (126) and the Silver–Müller condition (129), which implies that on $S_{\infty}$ the scattered field in Scenario II is $\mathcal{O}(r_{\infty}^{-1})$ and thus negligible compared with the unit-amplitude incident plane wave-the left-hand side of (130) reduces, to leading order to
$\frac{1}{\zeta_{0}}\frac{e^{ik_{\omega}r_{\infty}}}{r_{\infty}}\,\oint_{S_{\infty}}d^{2}\mathbf{r}\,\exp\left(ik_{\omega}r_{\infty}\mathbf{n}_{\infty}\cdot\mathbf{n}\right)\\ \left(\mathbf{A}_{i}(\mathbf{n})\times\mathbf{n}_{\infty}\times\mathbf{u}_{\infty}\cdot\mathbf{n}-\mathbf{u}_{\infty}\times\mathbf{n}\times\mathbf{A}_{i}(\mathbf{n})\cdot\mathbf{n}\right)$ (131)
We now use $d^{2}\mathbf{r}=r_{\infty}^{2}\sin\theta\,d\theta\,d\phi$:
$\frac{1}{\zeta_{0}}{e^{ik_{\omega}r_{\infty}}}\,\int_{0}^{2\pi}d\phi\int_{0}^{\pi}d\theta\,{r_{\infty}}\sin\theta\,\exp\left(ik_{\omega}r_{\infty}\mathbf{n}_{\infty}\cdot\mathbf{n}\right)\\ \left[\mathbf{A}_{i}(\mathbf{n})\times\left(\mathbf{n}_{\infty}\times\mathbf{u}_{\infty}\right)\cdot\mathbf{n}+\mathbf{u}_{\infty}\cdot\mathbf{A}_{i}(\mathbf{n})\right]$ (132)
Without loss of generality, choose $\mathbf{u}_{\infty}=\mathbf{x}$ and $\mathbf{n}_{\infty}=\mathbf{z}$. With $\mathbf{n}=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)$ and $\mathbf{A}_{i}(\mathbf{n})\times\left(\mathbf{n}_{\infty}\times\mathbf{u}_{\infty}\right)\cdot\mathbf{n}=\mathbf{x}\cdot\mathbf{A}_{i}\cos\theta-\mathbf{z}\cdot\mathbf{A}_{i}\sin\theta\cos\phi$, Eq. 132 becomes
$\frac{1}{\zeta_{0}}{e^{ik_{\omega}r_{\infty}}}\,\int_{0}^{2\pi}d\phi\int_{0}^{\pi}d\theta\,{r_{\infty}}\sin\theta\,\exp\left(ik_{\omega}r_{\infty}\cos\theta\right)\\ \left[\mathbf{A}_{i}(\theta,\phi)\cdot\mathbf{x}(1+\cos\theta)-\mathbf{A}_{i}(\theta,\phi)\cdot\mathbf{z}\sin\theta\cos\phi\right]$ (133)
The expression 133 contains integrals of the type
$\int_{-1}^{1}d\xi\,e^{i{k}_{\omega}r_{\infty}\xi}f(\xi),$
where $\xi=\cos\theta$, which can be integrated by parts to yield
$\frac{e^{ik_{\omega}r_{\infty}}f(1)-e^{-ik_{\omega}r_{\infty}}f(-1)}{ik_{\omega}r_{\infty}}+\mathcal{O}\left(\frac{1}{k_{\omega}^{2}r_{\infty}^{2}}\right),$
provided that $df/d\xi$ is bounded. Keeping the leading term, eq. 133 gives
$\frac{1}{\zeta_{0}}r_{\infty}{e^{ik_{\omega}r_{\infty}}}{2\pi}\frac{2e^{ik_{\omega}r_{\infty}}}{ik_{\omega}r_{\infty}}\left.\mathbf{A}_{i}\cdot\mathbf{x}\right|_{\theta=0}\\ =-i\frac{4\pi}{\zeta_{0}}\frac{1}{k_{\omega}}e^{2ik_{\omega}r_{\infty}}\left.\mathbf{A}_{i}\cdot\mathbf{x}\right|_{\theta=0}$ (134)
Plugging back this expression in Eq. 130 and using $1/\zeta_{0}=\varepsilon_{0}c$ and $\omega=ck_{\omega}$ gives
$\left.\mathbf{A}_{i}\cdot\mathbf{x}\right|_{\theta=0}=\frac{k_{\omega}^{2}}{4\pi\varepsilon_{0}}\mu_{i}\,\mathbf{F}_{\omega\mathbf{z}\mathbf{x}}(\mathbf{r}_{i})\cdot\mathbf{u}_{i}$ (135)
Going back to the original reference system, we obtain the projection identity
$\mathbf{A}_{i}(\mathbf{n})\cdot\mathbf{u}_{\infty}=\frac{k_{\omega}^{2}}{4\pi\varepsilon_{0}}\,\mu_{i}\,\mathbf{F}_{\omega\mathbf{n}\mathbf{u}_{\infty}}(\mathbf{r}_{i})\cdot\mathbf{u}_{i},$ (136)
up to the overall phase factor that cancels in the power observables. In the scattering-mode normalization introduced in Eq. (9), Eq. 136 becomes:
$\mathbf{A}_{i}(\mathbf{n})\cdot\mathbf{u}_{\infty}=\sqrt{\frac{\pi\omega}{\hbar\varepsilon_{0}c}}\,\mu_{i}\,[\mathbf{E}_{\omega\mathbf{n}\mathbf{u}_{\infty}}(\mathbf{r}_{i})\cdot\mathbf{u}_{i}].$ (137)
Relation  (137) states that the $\mathbf{u}_{\infty}$-polarized far-field amplitude in direction $\mathbf{n}$ is proportional to the projection of the normalized scattering mode $\mathbf{E}_{\omega\mathbf{n}\bm{\nu}}$ at the dipole location onto the dipole orientation $\mathbf{u}_{i}$. In other words, it links the local coupling between the dipole and the scattering mode at $\mathbf{r}_{i}$ to the corresponding far-field pattern produced when the object is driven by a dipole $\mu_{i}\mathbf{u}_{i}$ at $\mathbf{r}_{i}$.